### v1

# DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures

Abstract
Large language models (LLMs) are increasingly deployed for real-world tasks that fundamentally involve data manipulation. A core requirement across these tasks is the ability to perform structural reasoning—that is, to understand and reason about data relationships. For example, customer requests require a temporal ordering, which can be represented by data structures such as queues. However, existing benchmarks primarily focus on high-level, application-driven evaluations without isolating this fundamental capability. To address this gap, we introduce DSR-Bench, a novel benchmark evaluating LLMs’ structural reasoning capabilities through data structures, which provide interpretable representations of data relationships. DSR-Bench includes 20 data structures, 35 operations, and 4,140 problem instances, organized hierarchically for fine-grained analysis of reasoning limitations. Our evaluation pipeline is fully automated and deterministic, eliminating subjective human or model-based judgments. Its synthetic nature also ensures scalability and minimizes data contamination risks. We benchmark nine state-of-the-art LLMs. Our analysis shows that instruction-tuned models struggle with basic multi-attribute and multi-hop reasoning. Furthermore, while reasoning-oriented models perform better, they remain fragile on complex and hybrid structures, with the best model achieving an average score of only 47% on the challenge subset. Crucially, models often perform poorly on multi-dimensional data and natural language task descriptions, highlighting a critical gap for real-world deployment.
1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across practical tasks such as database management [62], coding [3, 58], and scheduling [20]. These tasks differ in domain, but they share a core requirement: the ability to understand and manipulate data according to relationships such as order, hierarchy, and connectivity—a capability we refer to as structural reasoning.
Structural reasoning is an essential, yet often implicit, requirement in problem solving. For instance, when using an LLM to plan a trip, the system may need to interpret a map of cities as a graph, construct an itinerary by finding a path through that graph, and prioritize visits using a queue. As LLMs are increasingly deployed in real-world applications, particularly safety-critical ones such as robotics [39], healthcare [38], and autonomous systems [14, 1], it becomes imperative to rigorously evaluate their structural reasoning capabilities to identify potential failure modes prior to deployment. Furthermore, LLMs are known to struggle with complex forms of reasoning, such as spatial reasoning [22, 55], underscoring the timely need for a systematic evaluation of their structural reasoning abilities.
However, existing benchmarks offer limited insight into structural reasoning. Most evaluations focus on high-level, domain-specific tasks such as mathematical problem solving [28, 25, 50], competitive programming [16], or STEM-related questions [15]. These tasks are typically multi-step and complex, making it difficult to isolate structural reasoning performance from other reasoning abilities. As a result, a critical gap remains: there are no targeted, fine-grained benchmarks for evaluating structural reasoning, which is a core skill underlying many downstream applications
To address this gap, we propose evaluating structural reasoning through the lens of data structures. Data structures offer a controlled and interpretable framework for modeling diverse data relationships: arrays for sequences, stacks and queues for temporal logic, trees for hierarchies, and graphs for complex networks. Since structural reasoning manifests as operations on these canonical structures, data structure tasks provide a principled way to probe LLM reasoning abilities at a granular level. Such a fine-grained evaluation can not only identify reliable applications of LLMs for safe deployment, but also pinpoint areas where reasoning breaks down and how it might be improved.
Building on this insight, we introduce DSR-Bench, a benchmark designed to isolate and evaluate LLM structural reasoning via data structure manipulation tasks. DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories, see Figure 1) and 35 operations across three length types (short, medium, and long). We also introduce DSR-Bench-challenge, a subset of DSR-Bench consisting of data structures with particularly complex designs. Each task probes whether the model can construct, maintain, and reason about data relationships within an interpretable and automatically verifiable setting.
DSR-Bench offers several key advantages for evaluating structural reasoning:
Hierarchical organization. Tasks are organized by increasing structural complexity, enabling a fine-grained analysis of specific reasoning skills. Within each category, we design a range of tasks to isolate different sources of complexity, allowing structural reasoning to be broken down into progressively more challenging tasks. This approach precisely identifies the specific types of data relationships that pose difficulties for LLMs.
Deterministic evaluation. Each data structure task has a concise and well-defined final state as its ground-truth label, allowing for deterministic and unambiguous scoring. Unlike open-ended tasks, this design supports fully automated evaluation without the need for human or model-based judgment, resulting in a fairer and more objective evaluation pipeline [4, 10, 54].
We benchmark nine state-of-the-art LLMs, including both instruction-tuned and reasoning models. Our analysis reveals the following key insights:
1.
Instruction-tuned models struggle with multi-attribute and multi-hop reasoning. Their accuracy drops sharply on tasks involving multi-attribute data (such as hashmaps) and tasks requiring the maintenance of multi-hop properties (such as red-black trees). Prompting strategies that encourage intermediate steps can improve performance, but techniques like CoT tend to be most effective when the data structure is less standard.
2.
Reasoning models still struggle with complex structures and non-standard tasks. Despite outperforming instruction-tuned models, reasoning models perform poorly on structurally complex tasks, with the best scoring only 47% on DSR-Bench-challenge. Some also fail when task instructions conflict with what are likely learned assumptions (e.g., consistently predicting tree depth as the number of edges instead of nodes, despite explicit instructions). These errors suggest a reliance on entrenched priors, limiting the model’s ability to adapt to user-defined constraints.
3.
Model performance declines significantly on complex spatial data structures. Accuracy drops as data dimensionality increases and degrades further when data is non-uniform. These patterns suggest a reliance on memorization over robust generalization or genuine reasoning.
4.
LLMs perform worse when data structure tasks are presented in natural language. Translating formal task descriptions into natural language leads to a marked drop in performance, posing a barrier to the practical deployment of LLMs for real-world problem solving.
We release all code at https://github.com/dransyhe/DSR-Bench and all data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 to support full reproducibility. We welcome community engagement and collaboration.
2 DSR-Bench: The Data Structure Reasoning Benchmark
In DSR-Bench, we propose a taxonomy of data structures grouped into six categories based on the types of data relationships they encode: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid. This taxonomy captures the diversity of relational patterns found in real-world data and enables systematic evaluation of LLM capabilities across distinct relationship types. Within each category, we include a spectrum of data structures that vary in complexity and operational constraints, allowing us to isolate specific reasoning challenges and precisely diagnose where models struggle.
2.1 Data structure tasks
DSR-Bench includes the following data structures and tasks. A detailed description of each data structure is provided in Section A.2.
Linear (Sequential): This category includes array and its operations: access, insert, delete, reverse, and search. Linear structures are fundamental as they introduce ordered relationships, enabling reasoning about position, sequence, and iteration, and they serve as a foundation for understanding more complex data abstractions.
Temporal (Time-based ordering): Temporal structures include stack, queue, priority queue, and skip list, which operate under last-in-first-out, first-in-first-out, priority-based, or probabilistic rules. Their operations are compound actions of insertions and deletions. Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers, and introduces an additional ordering constraint beyond simple linear sequencing.
Associative (Key-value mapping): We evaluate hashmap, trie, and suffix tree through tasks focused on construction and compound insertion-deletion operations. These tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures, which are core skills for efficient lookup and structured access in systems such as databases.
Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, with tasks involving traversals and compound operations. These structures are common in file systems and database indexing. These tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU). Graph tasks focus on breadth- and depth-first traversal, testing a model’s ability to reason about connectivity and explore many-to-many relationships, such as user interactions in social networks. DSU tasks centre on union-find operations, which are fundamental for dynamically managing group membership and identifying connected components in partitioned data.
Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple relational principles. This category includes LRU cache, which integrates temporal and memory constraints; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph. Tasks involve compound insertions and deletions, testing whether models can reason over combined constraints and generalize beyond individual structural concepts.
2.2 Data generation
We design a diverse set of operation tasks for each data structure, as detailed in Table 8. Tasks fall into three categories: construction, inspection (e.g., access, pre/in/post-order traversal), and manipulation (e.g., insert, delete). In addition to these atomic operations, which serve as fundamental building blocks and allow fine-grained diagnosis of specific failure modes, DSR-Bench also includes compound operations. These are composed of a sequence of manipulations (e.g., insert and delete) and are designed to test a model’s ability to maintain structural integrity across multiple steps.
All data is synthetically generated, which is efficient and less susceptible to training data contamination [56, 51]. Numerical inputs are sampled uniformly from the range 0 to 100, while string inputs are composed of lowercase English letters. For each task, ground-truth outputs are generated by implementing the corresponding data structure and executing the prescribed operations. The expected format of the final state is explicitly specified in the prompt. For simple structures like arrays, the final state is a sequence of elements. For more complex structures such as trees and graphs, the final state is represented using a canonical traversal, such as pre-order or breadth-first. The modular design of our benchmark also supports easy incorporation of new data structure types.
2.3 Suites
We categorize each task into three levels based on input length: short (5–10), medium (11–20), and long (21–30), to evaluate length generalization. For atomic operations, length corresponds to the number of input elements, such as list items or tree nodes. For compound operations, tasks start from an empty structure, and length is defined by the number of sequential operations applied.
We also curated a DSR-Bench-challenge subset highlighting data structures with particularly complex designs, aimed at pushing the limits of state-of-the-art models. To further assess structural reasoning capabilities, we also introduce two extension suites: DSR-Bench-spatial, focusing on spatial reasoning on multi-dimensional inputs (details see Section 4.2), and DSR-Bench-natural, evaluating reasoning over natural language descriptions (details see Section 4.3).
2.4 Prompt design
For each task, we design a prompt template and populate it with synthetic data to produce individual problem instances. Each prompt follows a consistent format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final computed outcome. Following recommended practices in prompt engineering [35], we also append the instruction “Answer the question in <token> tokens” to guide models toward producing concise outputs within a specified token budget. We also implement five different prompting methods for each task, as we detail in Section 3.2.
2.5 Structured output
A key advantage of our tasks is that model outputs can be deterministically verified using automated evaluation. This is possible because the final states of data structures can be represented as sequences of numbers or strings, enabling precise comparison. Unlike benchmarks involving proof writing or complex math problems, which often yield long and ambiguous outputs requiring subjective human or LLM-based judging [4, 10, 54], our evaluation framework is more objective and reproducible.
For automatic verification, models must produce outputs matching predefined JSON schemas enforced with Pydantic. For instance, a Binary Search Tree (BST) pre-order traversal returns a list[int]. We ensure that output formats are easy to parse: about 70% of the tasks yield a 1D integer array.
2.6 Scoring system
We evaluate each model by comparing its output to the ground-truth answer. Each correct output scores 1, otherwise 0. The overall performance is reported as the average accuracy across all categories and length levels. To ensure deterministic outputs and eliminate ambiguity during evaluation, every task admits a single, well-defined solution. Potential sources of nondeterminism, such as definitions of hash function and tie-breaking rules, are explicitly specified in the prompt and controlled during problem generation. Additionally, we run extensive tests across all tasks and manually inspect failure cases to verify that errors are not caused by ambiguity in the question statements or task specifications.
3 Experiment settings
In this section, we describe our experimental settings. Section 3.1 outlines the models evaluated, and Section 3.2 details the prompting methods used.
3.1 Models
We evaluate a broad set of LLMs across two categories: general-purpose instruction-tuned models and models designed for advanced reasoning. From each major provider, we select the latest publicly available flagship models. Each prompt is evaluated three times, and we report both the average and standard deviation, resulting in a total of 166,230 query evaluations.
Instruction-tuned LLMs are widely used in real-world applications due to their low latency, reduced cost, and scalability. Although not explicitly trained for algorithmic or mathematical reasoning, their ability to process structured data is critical in practical settings. We test GPT-4.1-2025-04-14 (OpenAI) [36], Llama3.3-70B (Meta) [13], Gemini-2.0-Flash-001 (Google DeepMind) [45] , Claude-3-5-Sonnet-20241022 (Anthropic) [2], and DeepSeek-V3 (DeepSeek) [7].
We also evaluate a suite of models optimized for advanced reasoning. These models are explicitly trained to handle complex, multi-step tasks by maintaining intermediate states, including: o4-mini-2025-04-16 (OpenAI) [34], Gemini-2.5-Pro-Preview-03-25 (Google DeepMind) [45], Claude-3-7-Sonnet-20250219 (Anthropic) [2], and DeepSeek-R1 (DeepSeek) [7].
3.2 Prompts
To examine how prompting strategies affect model performance on data structure tasks, we evaluate five configurations. Unlike reasoning models with internal mechanisms for multi-step inference (e.g., explicit reasoning tokens), instruction-tuned models are especially sensitive to prompt formulation. Our prompting configurations include: (i) Stepwise, which adds a “steps” field to the JSON output; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single reasoning example with a final output; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text. See Section A.3 for examples.
4 Evaluation
In this section, we present experimental results evaluating all nine LLMs and highlight their strengths and weaknesses across structural reasoning tasks on DSR-Bench and DSR-Bench-challenge (Section 4.1), DSR-Bench-spatial (Section 4.2), and DSR-Bench-natural (Section 4.3).
4.1 Can LLMs understand and manipulate data structures?
In this section, we present our experimental results of LLMs on DSR-Bench (including DSR-Bench-challenge). In Section 4.1.1, we discuss the results of the five instruction-tuned models, and in Section 4.1.2, we highlight key observations on the four reasoning models.
| Relationship | Data Structure | o4- mini | Gemini- 2.5-Pro | Claude-3. 7-Sonnet | DeepSe ek-R1 | GPT- 4.1 | Gemini- 2.0-Flash | Claude-3. 5-Sonnet | DeepSe ek-V3 | Llama 3.3-70B |
| Linear | Array | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 |
| Category avg. | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 | |
| Temporal | Stack | 1.00 | 1.00 | 0.97 | 0.99 | 0.55 | 0.36 | 0.99 | 0.41 | 0.04 |
| Queue | 1.00 | 1.00 | 1.00 | 0.98 | 0.55 | 0.36 | 0.99 | 0.43 | 0.25 | |
| LRU | 1.00 | 1.00 | 0.30 | 0.16 | 0.85 | 0.78 | 0.82 | 0.01 | 0.50 | |
| Priority Queue | 0.55 | 0.51 | 0.28 | 0.65 | 0.25 | 0.16 | 0.33 | 0.20 | 0.08 | |
| Category avg. | 0.89 | 0.84 | 0.64 | 0.69 | 0.55 | 0.42 | 0.79 | 0.26 | 0.22 | |
| Associative | Hashmap | 0.51 | 0.64 | 0.63 | 0.33 | 0.06 | 0.10 | 0.16 | 0.00 | 0.00 |
| Trie | 0.68 | 0.80 | 0.08 | 0.49 | 0.18 | 0.17 | 0.49 | 0.02 | 0.00 | |
| Suffix Tree | 0.73 | 0.85 | 0.91 | 0.96 | 0.00 | 0.01 | 0.08 | 0.67 | 0.00 | |
| Skip List | 0.62 | 0.73 | 0.75 | 0.69 | 0.07 | 0.06 | 0.42 | 0.02 | 0.01 | |
| Category avg. | 0.63 | 0.76 | 0.59 | 0.62 | 0.08 | 0.09 | 0.29 | 0.18 | 0.00 | |
| Hierarchical | BST | 0.86 | 0.86 | 0.64 | 0.73 | 0.59 | 0.43 | 0.71 | 0.58 | 0.34 |
| Heap | 0.68 | 0.61 | 0.40 | 0.48 | 0.20 | 0.10 | 0.27 | 0.15 | 0.07 | |
| RB tree | 0.65 | 0.77 | 0.30 | 0.62 | 0.12 | 0.12 | 0.46 | 0.09 | 0.05 | |
| B+ tree | 0.97 | 0.95 | 0.38 | 0.31 | 0.23 | 0.12 | 0.23 | 0.08 | 0.01 | |
| K-D Tree | 0.47 | 0.71 | 0.00 | 0.45 | 0.00 | 0.01 | 0.03 | 0.00 | 0.00 | |
| K-D Heap | 0.10 | 0.05 | 0.05 | 0.11 | 0.04 | 0.04 | 0.05 | 0.03 | 0.01 | |
| Category avg. | 0.64 | 0.78 | 0.33 | 0.45 | 0.20 | 0.14 | 0.34 | 0.16 | 0.08 | |
| Network | Graph | 0.87 | 0.94 | 0.11 | 0.67 | 0.15 | 0.05 | 0.06 | 0.06 | 0.02 |
| DSU | 0.98 | 1.00 | 0.99 | 1.00 | 0.02 | 0.00 | 0.02 | 0.82 | 0.00 | |
| Geom Graph | 0.16 | 0.13 | 0.02 | 0.13 | 0.07 | 0.02 | 0.02 | 0.01 | 0.02 | |
| Category avg. | 0.67 | 0.69 | 0.38 | 0.60 | 0.08 | 0.03 | 0.03 | 0.30 | 0.01 | |
| Hybrid | Bloom Filter | 0.61 | 0.89 | 0.16 | 0.74 | 0.04 | 0.04 | 0.04 | 0.04 | 0.02 |
| DAWG | 0.25 | 0.07 | 0.06 | 0.14 | 0.05 | 0.06 | 0.07 | 0.06 | 0.01 | |
| Category avg. | 0.43 | 0.48 | 0.11 | 0.44 | 0.05 | 0.05 | 0.06 | 0.05 | 0.02 | |
| Score | Overall avg. | 0.72 | 0.76 | 0.51 | 0.63 | 0.31 | 0.27 | 0.41 | 0.32 | 0.17 |
4.1.1 Instruction-tuned models
Instruction-tuned models struggle with multi-attribute reasoning.
As shown in Table 1, these models show sharp accuracy drops in tasks involve elements with multiple attributes. For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority. Similarly, in the hashmap task, manual inspection of errors shows that models confuse keys and values, delete the wrong items, or hallucinate entries. These results reveal a key limitation for real-world deployment, where managing entities with multiple interacting properties, such as deadlines, priorities, or key-value records, is common.
Multi-hop reasoning in hierarchical or network structures remains a key challenge.
We see from Table 1 that, while models perform reliably on binary search trees, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across multiple ancestral levels. Performance declines further on B+ trees, which require reasoning over wider spans of child pointers, and on graph traversal tasks involving many-to-many relationships. These results underscore a persistent limitation: instruction-tuned LLMs struggle to maintain and manipulate context beyond local neighborhoods.
| Model | Stepwise | 0-CoT | CoT | 3-shot | None |
| DeepSeek-V3 | 0.67 | 0.67 | 0.66 | 0.64 | 0.55 |
| Llama-3.3 | 0.46 | 0.46 | 0.48 | 0.46 | 0.34 |
| GPT-4.1 | 0.82 | 0.83 | 0.94 | 0.80 | 0.59 |
| Gemini-2.0-Flash | 0.57 | 0.86 | 0.59 | 0.87 | 0.58 |
| Claude-3.5-Sonnet | 0.72 | 0.92 | 0.72 | 0.93 | 0.69 |
Prompting can help, but only when carefully designed.
As shown in Table 2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial. Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures. A representative case is suffix-tree construction: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section A.5).
4.1.2 Reasoning models
| Model | Score | Priority Queue | Suffix Tree | Trie | Skip List | Heap | Red-bla ck tree | B+ Tree | K-D Tree | K-D Heap | DSU | Geom Graph | Bloom Filter | DAWG |
| Gemini-2.5-Pro | 0.47 | 0.24 | 0.89 | 0.59 | 0.61 | 0.62 | 0.60 | 0.96 | 0.67 | 0.00 | 0.99 | 0.00 | 0.69 | 0.00 |
| DeepSeek-R1 | 0.36 | 0.48 | 0.90 | 0.05 | 0.54 | 0.27 | 0.37 | 0.21 | 0.01 | 0.01 | 0.99 | 0.01 | 0.31 | 0.00 |
| o4-mini | 0.34 | 0.30 | 0.37 | 0.32 | 0.41 | 0.71 | 0.37 | 0.94 | 0.38 | 0.00 | 0.94 | 0.00 | 0.07 | 0.06 |
| Claude-3.7-Sonnet | 0.21 | 0.04 | 0.79 | 0.00 | 0.61 | 0.13 | 0.12 | 0.13 | 0.00 | 0.00 | 0.98 | 0.00 | 0.00 | 0.00 |
Reasoning models remain brittle on complex and spatial data structures.
As shown in Table 1, reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures. However, overall score remains below 0.5 on DSR-Bench-challenge (Table 3), in particular for long tasks and complex data structures. For example, the highest score on heap is only 0.62, despite it being a widely used structure in scheduling and resource allocation. Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Appendix A.4). To further probe these limitations, we introduce DSR-Bench-spatial, described in Section 4.2.
Implicit priors may hinder instruction following.
We observe that while GPT-4.1 scores 0.66 on the binary search tree depth task (short), o4-mini performs much worse, with a score of just 0.01 (Section A.4). Manual inspection of incorrect outputs reveals that o4-mini consistently interprets depth as the number of edges rather than nodes, despite explicit instructions. Similarly, in an ablation on K-D Heap (Table 5), switching from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons. Directly quering o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys. These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors, limiting their reliability on tasks with strict user-defined constraints.
| K-D Heap | K-D Tree | Geom. Graph | ||||
| GPT-4.1 | o4-mini | GPT-4.1 | o4-mini | GPT-4.1 | o4-mini | |
| 1 | ||||||
| 2 | ||||||
| 3 | ||||||
| 5 | ||||||
| Lexicographic | Euclidean | |
| 1 | 0.79 | 0.82 |
| 2 | 0.87 | 0.34 |
| 3 | 0.84 | 0.26 |
| 5 | 0.90 | 0.21 |
4.2 Can LLMs reason structurally on spatial data?
Real-world data is often represented in multi-dimensional feature spaces. To assess whether LLMs can reason over such spatial data, we extend the benchmark with DSR-Bench-spatial, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space. These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics. Given task complexity, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps for Section 4.2 and Section 4.3.
Performance degrades as input dimensionality increases.
As shown in Table 5, accuracy consistently drops for both models as dimensionality increases. Higher-dimensional data challenges models to reason over more complex distance metrics and partitions, limiting their effectiveness in spatial tasks. For instance, K-D trees are commonly used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision [40]. However, a notable exception is the Geometric Graph, where 2D data performs best, likely because it is more commonly found in textbook examples used as training data.
| Distribution | GPT-4.1 | o4-mini |
| Uniform | ||
| Moon | ||
| Blob | ||
| Circle |
Limited robustness to non-uniform data distributions.
We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data. Specifically, we test K-D tree construction tasks using three non-uniform distributions from scikit-learn [37]: circles, moons, and blobs (examples shown in Figure 3, Section A.6). As shown in Table 6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in training data. Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning. In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts.
4.3 Can LLMs reason structurally on natural language?
While the previous sections evaluated LLMs on canonical data structures, real-world use-cases are often expressed in natural language. To bridge this gap, we extend the benchmark with DSR-Bench-natural, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
| Model | Task | Formal | Natural |
| GPT-4.1 | Queue | 1.00 | 0.77 |
| BST | 0.88 | 0.59 | |
| Graph | 0.42 | 0.43 | |
| o4-mini | Queue | 1.00 | 0.83 |
| BST | 1.00 | 0.93 | |
| Graph | 0.84 | 0.67 |
We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling). Synthetic data follows the same distributions as Section 2.2, with realistic substitutions (e.g., names for integers). Each scenario was written by humans, paraphrased by GPT-4o. All prompts were reviewed by three annotators for clarity and unambiguity. Details of prompt generation can be found in Section A.7.
LLMs struggle when shifting from formal to natural language.
As shown in Table 7, performance drops when tasks are described in natural language from formal descriptions, despite identical operation distributions. This suggests that even reasoning models struggle to apply their abilities in language-rich, real-world contexts. Bridging this gap is crucial for reliable deployment and presents a key direction for future research. Furthermore, the higher accuracy on formal descriptors may stem from LLMs’ exposure to textbook-style patterns, where integers and explicit syntax are common.
5 Related Works
LLM reasoning benchmarks
Many benchmarks have been proposed to evaluate the capabilities and limitations of LLMs (see Section A.1 for a detailed review). Reasoning-focused benchmarks have primarily been domain-specific, targeting areas such as math [5, 28, 25], coding challenges [50, 16, 3, 58], STEM question answering [15], and logic puzzles [50, 11]. These often require long, complex responses involving intertwined reasoning steps, and rely on human or LLM-based judgment for evaluation [4, 10, 54]. In contrast, we take a more foundational approach: testing whether LLMs can understand and manipulate data relationships, which is a core capability that underlies reasoning in math and programming, isolating structural reasoning from other complexities.
Algorithmic reasoning with LLMs
Prior work on algorithmic reasoning has primarily focused on arithmetic tasks and length generalization in transformers [59, 60, 21], or on specific graph problems such as cycle detection and connectivity [48, 8]. These efforts offer limited coverage of broader data structure reasoning. One closely related work, CLRS-Text [32], evaluates whether LLMs can simulate 30 algorithms from the CLRS benchmark suite [47, 6] via textual inputs. Rather than testing whether LLMs can replicate specific algorithms, we evaluate their ability to reason over data structures, which are core abstractions and prerequisites for algorithmic execution. This offers a more general and interpretable assessment of algorithmic reasoning, revealing not only whether models succeed, but also which types of structural reasoning they lack.
6 Conclusion, limitations, and future work
We introduce DSR-Bench, the first benchmark for evaluating LLMs’ structural reasoning through canonical data structures. DSR-Bench enables fine-grained, controlled, and objective evaluation using synthetic data and an automated scoring pipeline. Benchmarking nine state-of-the-art LLMs reveals key limitations across both instruction-tuned and reasoning models. Notably, DSR-Bench-challenge remains difficult even for the strongest models. In addition, we propose two extensions, DSR-Bench-spatial and DSR-Bench-natural, to evaluate spatial reasoning and natural language understanding of data structures. These results uncover further gaps, suggesting current LLMs are not yet ready for structure-dependent, real-world tasks.
Limitation and future work
While our scoring system compares the final state of the data structure, additional insights could be gained by evaluating models’ intermediate thought traces. However, such traces are unavailable from some providers and can be costly to retrieve. As interfaces grow more transparent, we hope to incorporate such signals to enrich error analysis and better understand model reasoning. Additionally, while DSR-Bench covers a wide range of structures, expanding it to cover emerging representations, such as graph-probabilistic hybrid structures, would further broaden its scope. Our benchmark is modular and extensible, and we welcome community contributions to advance these directions and help develop an even more comprehensive evaluation framework.
Acknowledgments and Disclosure of Funding
We thank Han Xuanyuan, Connor Lawless, and members of the Vitercik Lab for their meaningful discussions and valuable feedback on the manuscript. This work was supported in part by NSF grant CCF-2338226.
Appendix A Appendix
A.1 Additional related works on LLM benchmarking
Large language models (LLMs) have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations. Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding [49, 9], text generation [33, 41], reasoning [9, 46], and machine translation [53]. Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness [42, 31, 19, 24, 57, 29]. Specialized benchmarks have also emerged in scientific and technical domains, including mathematics [50, 25], programming [50, 16, 3, 58], data analysis [43, 27, 50, 44, 17], and medicine [23, 26, 30, 52, 18].
While several studies examine how LLMs convert unstructured inputs into tabular or relational formats [44, 17], our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs? To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
A.2 Details of data structures and operations
In this section, we list the data structures and the corresponding operations tested in Table 8. We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
| Category | Data Structure | Description | Operation | Application |
| Linear | Array | Contiguous memory | Access, Delete, Insert, Reverse, Search | Data storage |
| Temporal | Stack | LIFO (Last-In, First-Out) | Compound (Push, Pop) | Syntax parsing |
| Queue | FIFO (First-In, First-Out) | Compound | OS management | |
| LRU Cache | Least-recently-used | Cache (Evict, Add) | Web broswers | |
| Priority Queue* | Priority ordering | Compound | Job scheduling | |
| Associative | Hashmap | Key-value storage | Compound | Large-scaled storage |
| Trie* | Hierarchical mapping of strings | Compound | Autocomplete | |
| Suffix Tree* | Text indexing via suffixes | Compound | DNA pattern matching | |
| Skip List* | Probabilistic layers for fast search | Compound | Concurrent databases | |
| Hierarchical | Binary Search Tree | Hierarchical storage | Pre/In/Post-Order Traversal, Insert, Remove, Compound | Computer networks |
| Heap* | Complete binary tree with priority ordering | Heapify, Compound | Memory management | |
| Red-Black Tree* | Self-balanced tree | Construct, Compound | Database indexes | |
| B+ Tree* | Multi-way balanced tree | Compound | File systems | |
| K-D Tree* | Hierarchical, spatial partition | Construct, Compound | 3D graphics | |
| K-D Heap* | Hierarchical, complete binary tree, high dimensional priority | Compound | GPU job scheduling | |
| Network | Graph | Many-to-many relationships | Breadth-First Traversal, Depth-First Traversal | Social networks |
| Disjoint Set Union* | Sets partition & union | Compound (Union, Find) | Physics simulation | |
| Geometric Graph* | Graph modeling spatial data | Construct | Public transportation | |
| Hybrid | Bloom Filter* | Probabilistic set and hashmap | Compound | Spam detection |
| Directed Acyclic Word Graph* | Graph and trie tree | Compound | Compilers | |
Array
An array contains a list of elements stored in contiguous memory. We test its access, deletion, reversal, and search operations. To remove ambiguity, we specify that the array is 0-indexed. For operations like deletion, if duplicates exist, we delete the first occurrence. The final state is a list of elements in the array.
Stack
A stack is a linear data structure that follows a Last-In-First-Out (LIFO) order. We test compound operations consisting of random sequences of push and pop from the top of the stack. The final state is a list of remaining elements in the stack.
Queue
A queue is a linear data structure that follows a First-In-First-Out (FIFO) order. We evaluate compound operations of enqueue to the back and dequeue from the front. The final state is a list of remaining elements in the queue.
LRU Cache
An LRU (Least Recently Used) cache stores a fixed number of items and evicts the least recently accessed one when full. We evaluate this caching operation with a sequence of requests as input. The final state is a set of elements in the LRU cache.
Priority Queue
A priority queue stores elements with integer priorities, allowing access to the highest-priority element. We test compound operations including insert, remove, raise key, and decrease key, using a Fibonacci heap. Ties are broken by insertion order. The final state is a level-order traversal of the Fibonacci heap forest, outputting (value,priority) pairs, with nodes at each level sorted by descending priority and ties broken by larger value first.
Hashmap
A hashmap is a key-value structure supporting fast access, insertion, and deletion via hashed keys. We test compound insert and delete operations, specifying the hash function and using chaining for collision resolution. The final state is a list of key-value pairs per bucket, preserving insertion order within each chain.
Trie
A trie is a tree-based data structure for storing strings, where each node represents a character and paths from the root to leaves represent complete words, where common prefixes are shared. When generating strings, we increase the likelihood of shared prefixes to ensure the resulting trie has meaningful structure. The final state is a pre-order traversal of the trie, where each node’s children are visited in lexicographical order to ensure an unambiguous representation.
Suffix Tree
A suffix tree is a compressed trie built from all suffixes of a string, where each edge can represent multiple characters and each path from the root corresponds to a substring. We test the construction of a suffix tree from a given word, appending a terminal character “$” to ensure a unique structure. The final state is a pre-order traversal collecting edge labels, with child edges visited in lexicographical order and “$” taking priority.
Skip List
A skip list is a probabilistic data structure composed of multiple layers of linked lists, where higher layers allow “skipping” over elements for faster access. We test compound operations of insert and delete. Insertion begins at the bottom layer, with the element randomly promoted to higher levels; pointers are updated at each level to preserve the structure. To remove ambiguity, promotion probabilities are explicitly specified in the prompts. The final state is represented as a list of lists, each corresponding to a layer of the skip list.
Binary Search Tree (BST)
A binary search tree is a hierarchical structure where each node has at most two children: the left holds smaller values, and the right holds larger ones. We test insert, remove, tree traversals (pre-order, in-order, post-order), depth computation, and compound insert-remove operations. Inputs are guaranteed to contain no duplicates to ensure unique outputs. The final state for traversal tasks is a list of elements in the specified order, while for insert, remove, and compound tasks, it includes both pre-order and post-order traversals.
Heap
A heap is a complete binary tree that satisfies the min-heap property, where each parent node is less than or equal to its children. We test both heapify and compound insert-delete operations using an array-based heap. Comparisons follow min-heap ordering, with ties broken by preferring the left child. The final state is the array representation of the heap.
Red-Black (RB) Tree
A red-black tree is a self-balancing binary search tree where each node is colored red or black and must satisfy specific balance rules: no two consecutive red nodes are allowed, and all root-to-leaf paths must have the same number of black nodes. We test both construction and compound (insert, delete) operations. The final state is a pre-order traversal of the nodes, represented as tuples (value, color).
B+ Tree
A B+ tree is a multi-way search tree used in databases and filesystems, where values are stored in leaf nodes and internal nodes serve as routing indexes. Leaf nodes are linked for efficient range queries. We specify splitting and merging rules to ensure unambiguous, balanced updates during compound insert and delete operations. The final state is a pre-order traversal of nodes, with keys in each node sorted in ascending order.
K-D Tree
A K-D (k-dimensional) tree recursively partitions space by alternating the splitting axis at each level. Each node represents a point and divides the space into two halves based on a chosen coordinate. It is commonly used for spatial indexing, range queries, and nearest neighbor search. We test the construction of K-D trees across different dimensionalities, specifying the axis splitting sequence and tie-breaking rules (e.g., median selection for even-sized splits) to ensure consistency. The final state is a pre-order traversal of the tree.
K-D Heap
A K-D heap maintains heap order based on a -dimensional priority with a comparison metric, enabling efficient access to extremal points in multidimensional datasets. We test compound operations of insert and delete across different dimensionalities. We specify an array-based heap implementation, with comparisons based on Euclidean distance and tie-breaking rules that prefer the left child in case of a tie. The final state is a list of vectors representing the contents of the min-heap.
Graph
A graph is a collection of nodes connected by edges, which can be directed or undirected, and is used to model networks, dependencies, and paths. We define graphs using edge list statements and test both breadth-first and depth-first traversals from a given source node, visiting neighbors in ascending order. Node values are unique to ensure consistent outputs. The final state is the list of nodes visited during the traversal.
Disjoint Set Union (DSU)
A disjoint set union maintains a partition of elements into disjoint subsets, supporting efficient merges and membership queries. Internally, it forms a forest where each node points to a representative root. We test two operations: a sequence of unions between subsets, followed by queries for each element’s representative. To ensure consistency, we specify that lower-rank roots are always attached to higher-rank ones. The final state lists the representative root of each input element in its original order.
Geometric (Geom) Graph
Geometric graphs are graphs with nodes embedded in geometric space, typically Euclidean, where edges are formed based on spatial relationships such as proximity. They are widely used in robotics, computer graphics, and sensor networks where spatial structure is essential. We compute the Euclidean distance between each pair of points and add an edge if the distance is below a given threshold, assigning the edge a weight equal to that distance. The final state is a breadth-first traversal from a specified source node, exploring all neighbors at each level before proceeding. We specify the order of search based on the edge weights.
Bloom Filter
A (counting) Bloom filter is a compact, probabilistic data structure for set membership testing, guaranteeing no false negatives and allowing a tunable false positive rate. It uses multiple hash functions to map each element to several positions in a counter array, incrementing or decrementing counts. We test on compound operations of insert and delete. We specify the hash functions used in the prompt to avoid ambiguity. The final state is the array of counters representing the Bloom filter.
Directed Acyclic Word Graph (DAWG)
A Directed Acyclic Word Graph (DAWG) is a compressed data structure for storing a set of words, sharing both prefixes and suffixes. Nodes indicate whether they mark the end of a word, and edges are labeled with characters. Unlike a trie, a DAWG merges equivalent subtrees to reduce redundancy, making it well-suited for large static dictionaries and lexicon lookups. We test compound operations of insert and delete, specifying that merging should occur at the final step along with the merging rules. To ensure a meaningful structure, we increase the likelihood of generating words with shared prefixes. The final state is a breadth-first traversal from the root (an empty string), where each node is recorded by the prefix it represents and whether it marks the end of a word.
A.3 Examples of prompting strategies
In this section, we illustrate prompting methods using compound operations of queue as an example.
Stepwise
This method expliclty adds a step attribute in the structured output, guiding the model to produce operations in a sequential and interpretable manner.
0-CoT
This method appends the phrase “Let’s think step by step” to the prompt to encourage reasoning without providing exemplars.
CoT
This strategy provides a single example that includes both intermediate reasoning steps and the final answer.
3-shot
This strategy provides three input-output examples to guide the model through pattern matching and demonstration.
None
This method adds the instruction “No additional text needed” to prompt concise, direct answers that fit within the token limit and conform to the structured output format.
A.4 Accuracy by task and length level across all models
In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level. Table 9 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels. Table 10 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite. For detailed per-model results across all tasks and length levels, see Table 11 (o4-mini), Table 12 (Gemini-2.5-Pro), Table 13 (Claude-3.7-Sonnet), Table 14 (DeepSeek-R1), Table 15 (GPT-4.1), Table 16 (Gemini-2.0-Flash), Table 17 (Claude-3.5-Sonnet), Table 18 (DeepSeek-V3), and Table 19 (Llama-3.3).
| Category | DS | Length | GPT-4.1 | Gemini-2.0-Flash | Claude-3.5-Sonnet | DeepSeek-V3 | Llama-3.3 |
| Linear | Array | Short | 0.98 | 0.98 | 1.00 | 1.00 | 0.89 |
| Medium | 0.95 | 0.92 | 0.96 | 0.97 | 0.70 | ||
| Long | 0.88 | 0.88 | 0.91 | 0.96 | 0.48 | ||
| Temporal | Queue | Short | 0.97 | 0.67 | 1.00 | 0.84 | 0.58 |
| Medium | 0.49 | 0.37 | 1.00 | 0.38 | 0.12 | ||
| Long | 0.18 | 0.03 | 0.98 | 0.07 | 0.06 | ||
| Stack | Short | 0.97 | 0.67 | 1.00 | 0.70 | 0.09 | |
| Medium | 0.49 | 0.37 | 1.00 | 0.49 | 0.04 | ||
| Long | 0.18 | 0.03 | 0.98 | 0.04 | 0.00 | ||
| Associative | Hashmap | Short | 0.19 | 0.28 | 0.37 | 0.00 | 0.00 |
| Medium | 0.00 | 0.01 | 0.10 | 0.00 | 0.00 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | ||
| Hierarchical | BST | Short | 0.82 | 0.63 | 0.89 | 0.76 | 0.46 |
| Medium | 0.56 | 0.37 | 0.66 | 0.55 | 0.31 | ||
| Long | 0.39 | 0.29 | 0.57 | 0.43 | 0.26 | ||
| Network | Graph | Short | 0.41 | 0.15 | 0.15 | 0.16 | 0.06 |
| Medium | 0.05 | 0.02 | 0.02 | 0.02 | 0.01 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Category | Data Structure | Length | o4-mini | Gemini-2.5-Pro | Claude-3.7-Sonnet | DeepSeek-R1 |
| Temporal | Priority Queue | Short | 0.89 | 0.87 | 0.70 | 0.92 |
| Medium | 0.47 | 0.43 | 0.11 | 0.54 | ||
| Long | 0.30 | 0.24 | 0.04 | 0.48 | ||
| Associative | Trie | Short | 0.99 | 1.00 | 0.23 | 0.93 |
| Medium | 0.73 | 0.82 | 0.00 | 0.50 | ||
| Long | 0.32 | 0.59 | 0.00 | 0.05 | ||
| Suffix Tree | Short | 0.96 | 0.74 | 0.96 | 1.00 | |
| Medium | 0.87 | 0.92 | 0.98 | 0.98 | ||
| Long | 0.37 | 0.89 | 0.79 | 0.90 | ||
| Skip List | Short | 0.84 | 0.93 | 0.87 | 0.89 | |
| Medium | 0.60 | 0.64 | 0.76 | 0.63 | ||
| Long | 0.41 | 0.61 | 0.61 | 0.54 | ||
| Hierarchical | Heap | Short | 0.61 | 0.67 | 0.70 | 0.63 |
| Medium | 0.72 | 0.55 | 0.37 | 0.53 | ||
| Long | 0.71 | 0.62 | 0.13 | 0.27 | ||
| Red Black Tree | Short | 0.92 | 0.97 | 0.69 | 0.86 | |
| Medium | 0.67 | 0.73 | 0.11 | 0.63 | ||
| Long | 0.37 | 0.60 | 0.12 | 0.37 | ||
| B+ Tree | Short | 0.99 | 0.97 | 0.70 | 0.49 | |
| Medium | 0.98 | 0.93 | 0.32 | 0.23 | ||
| Long | 0.94 | 0.96 | 0.13 | 0.21 | ||
| K-D Heap | Short | 0.22 | 0.07 | 0.11 | 0.23 | |
| Medium | 0.09 | 0.08 | 0.03 | 0.08 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.01 | ||
| K-D Tree | Short | 0.59 | 0.64 | 0.00 | 1.00 | |
| Medium | 0.43 | 0.81 | 0.00 | 0.34 | ||
| Long | 0.38 | 0.64 | 0.00 | 0.01 | ||
| Network | DSU | Short | 1.00 | 1.00 | 1.00 | 1.00 |
| Medium | 1.00 | 1.00 | 1.00 | 1.00 | ||
| Long | 0.94 | 0.99 | 0.98 | 0.99 | ||
| Hybrid | Bloom Filter | Short | 1.00 | 0.99 | 0.44 | 0.99 |
| Medium | 0.77 | 1.00 | 0.03 | 0.92 | ||
| Long | 0.07 | 0.69 | 0.00 | 0.31 | ||
| DAWG | Short | 0.49 | 0.07 | 0.17 | 0.40 | |
| Medium | 0.20 | 0.14 | 0.00 | 0.02 | ||
| Long | 0.06 | 0.00 | 0.00 | 0.00 | ||
| Geom Graph | Short | 0.37 | 0.30 | 0.07 | 0.36 | |
| Medium | 0.10 | 0.10 | 0.00 | 0.01 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.01 |
A.4.1 Performance of o4-mini
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Reverse | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Queue | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| LRU Cache | Cache | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| Priority Queue* | Compound | 0.89 (0.02) | 0.47 (0.06) | 0.30 (0.03) | |
| Associative | Hashmap | Compound | 0.89 (0.04) | 0.37 (0.00) | 0.26 (0.08) |
| Trie* | Compound | 0.99 (0.02) | 0.73 (0.06) | 0.32 (0.05) | |
| Suffix Tree* | Construct | 0.96 (0.02) | 0.87 (0.03) | 0.37 (0.07) | |
| Skip List* | Compound | 0.84 (0.02) | 0.60 (0.03) | 0.41 (0.02) | |
| Associative | BST | Insert | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) |
| Remove | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.03) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Depth | 0.01 (0.02) | 0.03 (0.03) | 0.00 (0.00) | ||
| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | ||
| Heap* | Compound | 0.77 (0.06) | 0.86 (0.07) | 0.74 (0.05) | |
| Heapify | 0.44 (0.05) | 0.58 (0.04) | 0.67 (0.03) | ||
| RB Tree* | Construct | 0.90 (0.03) | 0.32 (0.13) | 0.05 (0.02) | |
| Compound | 0.97 (0.00) | 0.64 (0.04) | 0.26 (0.02) | ||
| B+ Tree* | Compound | 0.99 (0.02) | 0.98 (0.04) | 0.94 (0.02) | |
| K-D Tree* | Construct | 0.59 (0.07) | 0.43 (0.06) | 0.38 (0.10) | |
| K-D Heap* | Compound | 0.22 (0.02) | 0.09 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.99 (0.02) | 0.97 (0.03) | 0.72 (0.14) |
| Depth-First Traversal | 1.00 (0.00) | 0.88 (0.02) | 0.64 (0.10) | ||
| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.94 (0.04) | |
| Geom Graph* | Construct | 0.37 (0.00) | 0.10 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 1.00 (0.00) | 0.77 (0.09) | 0.07 (0.00) |
| DAWG* | Compound | 0.49 (0.10) | 0.20 (0.09) | 0.06 (0.05) |
A.4.2 Performance of Gemini-2.5-Pro
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 1.00 (0.00) | 0.98 (0.02) | 1.00 (0.00) | ||
| Reverse | 1.00 (0.00) | 0.98 (0.02) | 1.00 (0.00) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Queue | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| LRU Cache | Cache | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| Priority Queue* | Compound | 0.87 (0.03) | 0.43 (0.03) | 0.24 (0.05) | |
| Associative | Hashmap | Compound | 0.90 (0.06) | 0.51 (0.04) | 0.50 (0.07) |
| Trie* | Compound | 1.00 (0.00) | 0.82 (0.05) | 0.59 (0.08) | |
| Suffix Tree* | Construct | 0.74 (0.08) | 0.92 (0.08) | 0.89 (0.02) | |
| Skip List* | Compound | 0.93 (0.03) | 0.64 (0.05) | 0.61 (0.02) | |
| Associative | BST | Insert | 0.99 (0.02) | 0.94 (0.05) | 0.96 (0.02) |
| Remove | 0.99 (0.02) | 0.84 (0.05) | 0.90 (0.03) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Depth | 0.22 (0.02) | 0.17 (0.07) | 0.18 (0.04) | ||
| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
| Heap* | Compound | 0.91 (0.04) | 0.72 (0.11) | 0.71 (0.05) | |
| Heapify | 0.43 (0.21) | 0.37 (0.03) | 0.53 (0.09) | ||
| RB Tree* | Construct | 0.83 (0.05) | 0.60 (0.03) | 0.28 (0.07) | |
| Compound | 0.91 (0.02) | 0.47 (0.06) | 0.18 (0.08) | ||
| B+ Tree* | Compound | 0.97 (0.06) | 0.93 (0.03) | 0.96 (0.05) | |
| K-D Tree* | Construct | 0.64 (0.11) | 0.81 (0.05) | 0.67 (0.12) | |
| K-D Heap* | Compound | 0.07 (0.00) | 0.08 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.98 (0.04) | 0.99 (0.02) | 0.86 (0.12) |
| Depth-First Traversal | 0.98 (0.02) | 0.94 (0.04) | 0.86 (0.05) | ||
| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | |
| Geom Graph* | Construct | 0.30 (0.03) | 0.10 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.99 (0.02) | 1.00 (0.00) | 0.69 (0.02) |
| DAWG* | Compound | 0.07 (0.02) | 0.14 (0.08) | 0.00 (0.00) |
A.4.3 Performance of Claude-3.7-Sonnet
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 1.00 (0.00) | 1.00 (0.00) | 0.96 (0.02) | ||
| Reverse | 1.00 (0.00) | 0.98 (0.02) | 0.97 (0.00) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Queue | Compound | 1.00 (0.00) | 0.93 (0.00) | 0.98 (0.02) | |
| LRU Cache | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | |
| Priority Queue* | Compound | 0.70 (0.06) | 0.11 (0.02) | 0.04 (0.02) | |
| Associative | Hashmap | Compound | 0.71 (0.02) | 0.16 (0.05) | 0.04 (0.05) |
| Trie* | Compound | 0.94 (0.02) | 0.64 (0.07) | 0.31 (0.10) | |
| Suffix Tree* | Construct | 0.23 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Skip List* | Compound | 0.87 (0.06) | 0.76 (0.05) | 0.61 (0.11) | |
| Associative | BST | Insert | 0.96 (0.04) | 0.98 (0.04) | 0.79 (0.02) |
| Remove | 0.94 (0.02) | 0.91 (0.02) | 0.92 (0.02) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Pre-Order Traversal | 1.00 (0.00) | 0.99 (0.02) | 1.00 (0.00) | ||
| Post-Order Traversal | 1.00 (0.00) | 0.74 (0.02) | 0.93 (0.03) | ||
| Depth | 0.24 (0.11) | 0.13 (0.03) | 0.02 (0.04) | ||
| Compound | 0.90 (0.03) | 0.21 (0.02) | 0.24 (0.05) | ||
| Heap* | Compound | 0.70 (0.03) | 0.32 (0.05) | 0.13 (0.03) | |
| Heapify | 0.89 (0.02) | 0.62 (0.08) | 0.26 (0.02) | ||
| RB Tree* | Construct | 0.19 (0.05) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.57 (0.03) | 0.03 (0.00) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.80 (0.00) | 0.18 (0.02) | 0.23 (0.03) | |
| K-D Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.11 (0.02) | 0.03 (0.00) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.40 (0.03) | 0.08 (0.02) | 0.01 (0.02) |
| Depth-First Traversal | 0.50 (0.03) | 0.11 (0.02) | 0.00 (0.00) | ||
| DSU* | Compound | 0.04 (0.02) | 0.12 (0.04) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.07 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.17 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
A.4.4 Performance of DeepSeek-R1
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 0.98 (0.02) | 1.00 (0.00) |
| Delete | 0.99 (0.02) | 1.00 (0.00) | 0.98 (0.02) | ||
| Insert | 0.98 (0.02) | 0.99 (0.02) | 0.99 (0.02) | ||
| Reverse | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.94 (0.07) |
| Queue | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.00) | |
| LRU Cache | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.01) | |
| Priority Queue* | Compound | 0.92 (0.02) | 0.54 (0.07) | 0.48 (0.05) | |
| Associative | Hashmap | Compound | 0.44 (0.05) | 0.01 (0.02) | 0.03 (0.03) |
| Trie* | Compound | 0.54 (0.24) | 0.32 (0.04) | 0.12 (0.06) | |
| Suffix Tree* | Construct | 0.93 (0.07) | 0.50 (0.07) | 0.05 (0.05) | |
| Skip List* | Compound | 0.89 (0.04) | 0.63 (0.03) | 0.54 (0.02) | |
| Associative | BST | Insert | 1.00 (0.00) | 0.98 (0.02) | 0.90 (0.03) |
| Remove | 0.98 (0.02) | 0.93 (0.03) | 0.88 (0.05) | ||
| In-Order Traversal | 1.00 (0.00) | 0.99 (0.02) | 1.00 (0.00) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.00) | ||
| Depth | 0.02 (0.04) | 0.06 (0.04) | 0.06 (0.05) | ||
| Compound | 0.97 (0.03) | 0.84 (0.08) | 0.65 (0.07) | ||
| Heap* | Compound | 0.49 (0.04) | 0.23 (0.07) | 0.21 (0.05) | |
| Heapify | 0.34 (0.02) | 0.16 (0.08) | 0.08 (0.06) | ||
| RB Tree* | Construct | 0.88 (0.02) | 0.10 (0.06) | 0.00 (0.00) | |
| Compound | 0.91 (0.04) | 0.37 (0.10) | 0.03 (0.03) | ||
| B+ Tree* | Compound | 0.81 (0.02) | 0.88 (0.04) | 0.70 (0.06) | |
| K-D Tree* | Construct | 1.00 (0.00) | 0.34 (0.05) | 0.01 (0.02) | |
| K-D Heap* | Compound | 0.23 (0.00) | 0.08 (0.02) | 0.01 (0.02) | |
| Network | Graph | Breadth-First Traversal | 0.92 (0.02) | 0.90 (0.06) | 0.46 (0.05) |
| Depth-First Traversal | 0.80 (0.09) | 0.58 (0.04) | 0.22 (0.02) | ||
| DSU* | Compound | 0.64 (0.56) | 0.92 (0.04) | 0.83 (0.07) | |
| Geom Graph* | Construct | 0.36 (0.02) | 0.01 (0.02) | 0.01 (0.02) | |
| Network | Bloom Filter* | Compound | 0.99 (0.02) | 0.92 (0.02) | 0.31 (0.02) |
| DAWG* | Compound | 0.40 (0.12) | 0.02 (0.02) | 0.00 (0.00) |
A.4.5 Performance of GPT-4.1
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 0.91 (0.08) | 0.79 (0.02) | 0.54 (0.02) | ||
| Reverse | 0.98 (0.02) | 0.97 (0.00) | 0.86 (0.02) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 0.97 (0.00) | 0.49 (0.02) | 0.18 (0.04) |
| Queue | Compound | 0.82 (0.04) | 0.59 (0.04) | 0.19 (0.07) | |
| LRU Cache | Cache | 0.94 (0.02) | 0.80 (0.00) | 0.81 (0.02) | |
| Priority Queue* | Compound | 0.63 (0.03) | 0.10 (0.00) | 0.03 (0.00) | |
| Associative | Hashmap | Compound | 0.19 (0.07) | 0.00 (0.00) | 0.00 (0.00) |
| Trie* | Compound | 0.39 (0.07) | 0.13 (0.03) | 0.01 (0.02) | |
| Suffix Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Skip List* | Compound | 0.21 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Associative | BST | Insert | 0.79 (0.04) | 0.50 (0.03) | 0.14 (0.02) |
| Remove | 0.78 (0.04) | 0.58 (0.02) | 0.36 (0.04) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.94 (0.02) | ||
| Pre-Order Traversal | 1.00 (0.00) | 0.97 (0.00) | 0.98 (0.02) | ||
| Post-Order Traversal | 0.82 (0.02) | 0.51 (0.04) | 0.23 (0.06) | ||
| Depth | 0.66 (0.05) | 0.08 (0.05) | 0.03 (0.00) | ||
| Compound | 0.69 (0.02) | 0.26 (0.02) | 0.03 (0.00) | ||
| Heap* | Compound | 0.58 (0.02) | 0.01 (0.02) | 0.00 (0.00) | |
| Heapify | 0.57 (0.03) | 0.04 (0.02) | 0.00 (0.00) | ||
| RB Tree* | Construct | 0.12 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.31 (0.04) | 0.02 (0.02) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.27 (0.00) | 0.30 (0.00) | 0.13 (0.00) | |
| K-D Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.31 (0.05) | 0.09 (0.02) | 0.00 (0.00) |
| Depth-First Traversal | 0.50 (0.03) | 0.00 (0.00) | 0.00 (0.00) | ||
| DSU* | Compound | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.20 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.16 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
A.4.6 Performance of Gemini-2.0-Flash
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 0.96 (0.02) | 0.87 (0.03) | 0.77 (0.03) | ||
| Insert | 0.99 (0.02) | 0.96 (0.02) | 1.00 (0.00) | ||
| Reverse | 0.96 (0.02) | 0.78 (0.02) | 0.64 (0.04) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.00) | ||
| Temporal | Stack | Compound | 0.67 (0.00) | 0.37 (0.00) | 0.03 (0.00) |
| Queue | Compound | 0.87 (0.00) | 0.33 (0.00) | 0.10 (0.00) | |
| LRU Cache | Cache | 0.93 (0.00) | 0.86 (0.02) | 0.56 (0.02) | |
| Priority Queue* | Compound | 0.38 (0.02) | 0.10 (0.00) | 0.01 (0.02) | |
| Associative | Hashmap | Compound | 0.28 (0.05) | 0.01 (0.02) | 0.00 (0.00) |
| Trie* | Compound | 0.31 (0.02) | 0.18 (0.02) | 0.03 (0.00) | |
| Suffix Tree* | Construct | 0.00 (0.00) | 0.02 (0.02) | 0.00 (0.00) | |
| Skip List* | Compound | 0.16 (0.02) | 0.00 (0.00) | 0.03 (0.00) | |
| Associative | BST | Insert | 0.31 (0.02) | 0.27 (0.03) | 0.06 (0.04) |
| Remove | 0.63 (0.09) | 0.33 (0.03) | 0.13 (0.03) | ||
| In-Order Traversal | 0.87 (0.00) | 0.66 (0.02) | 0.71 (0.04) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.93 (0.00) | ||
| Post-Order Traversal | 0.63 (0.00) | 0.17 (0.00) | 0.10 (0.00) | ||
| Depth | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) | ||
| Compound | 0.51 (0.05) | 0.12 (0.04) | 0.10 (0.00) | ||
| Heap* | Compound | 0.32 (0.05) | 0.03 (0.00) | 0.02 (0.02) | |
| Heapify | 0.23 (0.06) | 0.00 (0.00) | 0.00 (0.00) | ||
| RB Tree* | Construct | 0.08 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.43 (0.03) | 0.07 (0.00) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.17 (0.00) | 0.13 (0.03) | 0.06 (0.05) | |
| K-D Tree* | Construct | 0.02 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
| Depth-First Traversal | 0.19 (0.02) | 0.00 (0.00) | 0.00 (0.00) | ||
| DSU* | Compound | 0.01 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.18 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
A.4.7 Performance of Claude-3.5-Sonnet
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 0.93 (0.00) | ||
| Insert | 1.00 (0.00) | 0.90 (0.00) | 0.90 (0.00) | ||
| Reverse | 1.00 (0.00) | 0.88 (0.02) | 0.72 (0.05) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.04) |
| Queue | Compound | 0.87 (0.00) | 0.67 (0.03) | 0.79 (0.02) | |
| LRU Cache | Cache | 0.99 (0.02) | 0.90 (0.07) | 0.58 (0.13) | |
| Priority Queue* | Compound | 0.63 (0.00) | 0.27 (0.03) | 0.09 (0.02) | |
| Associative | Hashmap | Compound | 0.37 (0.03) | 0.10 (0.00) | 0.00 (0.00) |
| Trie* | Compound | 0.89 (0.04) | 0.50 (0.03) | 0.07 (0.00) | |
| Suffix Tree* | Construct | 0.21 (0.02) | 0.03 (0.00) | 0.00 (0.00) | |
| Skip List* | Compound | 0.77 (0.06) | 0.30 (0.07) | 0.20 (0.07) | |
| Associative | BST | Insert | 0.80 (0.06) | 0.50 (0.06) | 0.51 (0.05) |
| Remove | 0.96 (0.04) | 0.87 (0.00) | 0.77 (0.00) | ||
| In-Order Traversal | 0.97 (0.03) | 0.94 (0.02) | 0.94 (0.02) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
| Post-Order Traversal | 1.00 (0.00) | 0.69 (0.08) | 0.54 (0.02) | ||
| Depth | 0.76 (0.07) | 0.34 (0.10) | 0.16 (0.13) | ||
| Compound | 0.77 (0.07) | 0.28 (0.05) | 0.09 (0.02) | ||
| Heap* | Compound | 0.78 (0.04) | 0.13 (0.00) | 0.11 (0.02) | |
| Heapify | 0.53 (0.12) | 0.08 (0.04) | 0.00 (0.00) | ||
| RB Tree* | Construct | 0.13 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.44 (0.02) | 0.03 (0.00) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.40 (0.00) | 0.28 (0.08) | 0.02 (0.02) | |
| K-D Tree* | Construct | 0.09 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.13 (0.00) | 0.02 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.17 (0.03) | 0.02 (0.02) | 0.00 (0.00) |
| Depth-First Traversal | 0.13 (0.03) | 0.02 (0.02) | 0.00 (0.00) | ||
| DSU* | Compound | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.06 (0.04) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.20 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
A.4.8 Performance of DeepSeek-V3
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 0.97 (0.00) | 0.97 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Reverse | 1.00 (0.00) | 0.92 (0.02) | 0.92 (0.02) | ||
| Search | 1.00 (0.00) | 0.97 (0.00) | 0.93 (0.00) | ||
| Temporal | Stack | Compound | 0.70 (0.03) | 0.49 (0.02) | 0.04 (0.02) |
| Queue | Compound | 0.84 (0.02) | 0.38 (0.02) | 0.07 (0.03) | |
| LRU Cache | Compound | 0.94 (0.02) | 0.77 (0.06) | 0.76 (0.02) | |
| Priority Queue* | Compound | 0.53 (0.03) | 0.06 (0.04) | 0.00 (0.00) | |
| Associative | Hashmap | Compound | 0.04 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
| Trie* | Compound | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Suffix Tree* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Skip List* | Compound | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Associative | BST | Insert | 0.93 (0.03) | 0.62 (0.02) | 0.46 (0.05) |
| Remove | 0.84 (0.04) | 0.80 (0.03) | 0.66 (0.02) | ||
| In-Order Traversal | 0.97 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Post-Order Traversal | 0.82 (0.02) | 0.53 (0.03) | 0.20 (0.03) | ||
| Depth | 0.92 (0.02) | 0.43 (0.03) | 0.08 (0.04) | ||
| Compound | 0.68 (0.02) | 0.12 (0.04) | 0.06 (0.02) | ||
| Heap* | Compound | 0.23 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Heapify | 0.59 (0.02) | 0.06 (0.02) | 0.00 (0.00) | ||
| RB Tree* | Construct | 0.09 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.30 (0.03) | 0.00 (0.00) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.14 (0.05) | 0.10 (0.03) | 0.00 (0.00) | |
| K-D Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.07 (0.00) | 0.03 (0.00) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.29 (0.05) | 0.04 (0.02) | 0.00 (0.00) |
| Depth-First Traversal | 0.22 (0.02) | 0.03 (0.00) | 0.00 (0.00) | ||
| DSU* | Compound | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.04 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.17 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
A.4.9 Performance of Llama-3.3
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 0.56 (0.04) | 0.38 (0.02) |
| Delete | 0.81 (0.08) | 0.68 (0.04) | 0.44 (0.05) | ||
| Insert | 0.76 (0.04) | 0.78 (0.02) | 0.29 (0.07) | ||
| Reverse | 0.91 (0.02) | 0.56 (0.02) | 0.34 (0.07) | ||
| Search | 0.97 (0.00) | 0.90 (0.00) | 0.93 (0.00) | ||
| Temporal | Stack | Compound | 0.09 (0.02) | 0.04 (0.05) | 0.00 (0.00) |
| Queue | Compound | 0.58 (0.11) | 0.12 (0.02) | 0.06 (0.02) | |
| LRU Cache | Compound | 0.74 (0.08) | 0.44 (0.11) | 0.31 (0.11) | |
| Priority Queue* | Compound | 0.21 (0.05) | 0.01 (0.02) | 0.02 (0.02) | |
| Associative | Hashmap | Compound | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
| Trie* | Compound | 0.01 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Suffix Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Skip List* | Compound | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Associative | BST | Insert | 0.37 (0.00) | 0.14 (0.02) | 0.03 (0.00) |
| Remove | 0.49 (0.04) | 0.30 (0.03) | 0.14 (0.04) | ||
| In-Order Traversal | 0.60 (0.06) | 0.61 (0.08) | 0.61 (0.04) | ||
| Pre-Order Traversal | 0.86 (0.11) | 0.81 (0.08) | 0.78 (0.11) | ||
| Post-Order Traversal | 0.31 (0.04) | 0.04 (0.02) | 0.00 (0.00) | ||
| Depth | 0.33 (0.03) | 0.23 (0.00) | 0.28 (0.05) | ||
| Compound | 0.26 (0.02) | 0.01 (0.02) | 0.00 (0.00) | ||
| Heap* | Compound | 0.17 (0.03) | 0.03 (0.00) | 0.00 (0.00) | |
| Heapify | 0.24 (0.05) | 0.00 (0.00) | 0.00 (0.00) | ||
| RB Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Compound | 0.31 (0.02) | 0.00 (0.00) | 0.00 (0.00) | ||
| B+ Tree* | Compound | 0.02 (0.04) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Tree* | Construct | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| K-D Heap* | Compound | 0.02 (0.02) | 0.02 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.07 (0.06) | 0.00 (0.00) | 0.00 (0.00) |
| Depth-First Traversal | 0.04 (0.05) | 0.01 (0.02) | 0.00 (0.00) | ||
| DSU* | Compound | 0.00 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Geom Graph* | Construct | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.00 (0.00) | 0.07 (0.00) | 0.00 (0.00) |
| DAWG* | Compound | 0.02 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
A.5 Accuracy by prompting methods across instruction-tuned models
This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None. The results are shown in Table 20 (GPT-4.1), Table 21 (Gemini-2.0-Flash), Table 22 (Claude-3.5-Sonnet), Table 23 (DeepSeek-V3), and Table 24 (Llama-3.3).
| Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
| Array | Access | |||||
| Delete | ||||||
| Insert | ||||||
| Reverse | ||||||
| Search | ||||||
| Queue | Compound | |||||
| Stack | Compound | |||||
| LRU Cache | Cache | |||||
| Priority Queue | Compound | |||||
| Hashmap | Compound | |||||
| Trie | Compound | |||||
| Suffix Tree | Construct | |||||
| Skip List | Compound | |||||
| BST | Insert | |||||
| Remove | ||||||
| In-Order Traversal | ||||||
| Pre-Order Traversal | ||||||
| Post-Order Traversal | ||||||
| Depth | ||||||
| Compound | ||||||
| Heap | Compound | |||||
| Heapify | ||||||
| RB Tree | Construct | |||||
| Compound | ||||||
| B+ Tree | Compound | |||||
| Graph | Breadth-First Traversal | |||||
| Depth-First Traversal | ||||||
| DSU | Compound | |||||
| Bloom Filter | Compound | |||||
| DAWG | Compound |
| Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
| Array | Access | |||||
| Delete | ||||||
| Insert | ||||||
| Reverse | ||||||
| Search | ||||||
| Queue | Compound | |||||
| Stack | Compound | |||||
| LRU Cache | Cache | |||||
| Priority Queue | Compound | |||||
| Hashmap | Compound | |||||
| Trie | Compound | |||||
| Suffix Tree | Construct | |||||
| Skip List | Compound | |||||
| BST | Insert | |||||
| Remove | ||||||
| In-Order Traversal | ||||||
| Pre-Order Traversal | ||||||
| Post-Order Traversal | ||||||
| Depth | ||||||
| Compound | ||||||
| Heap | Compound | |||||
| Heapify | ||||||
| RB Tree | Construct | |||||
| Compound | ||||||
| B+ Tree | Compound | |||||
| Graph | Breadth-First Traversal | |||||
| Depth-First Traversal | ||||||
| DSU | Compound | |||||
| Bloom Filter | Compound | |||||
| DAWG | Compound |
| Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
| Array | Access | |||||
| Delete | ||||||
| Insert | ||||||
| Reverse | ||||||
| Search | ||||||
| Queue | Compound | |||||
| Stack | Compound | |||||
| LRU Cache | Cache | |||||
| Priority Queue | Compound | |||||
| Hashmap | Compound | |||||
| Trie | Compound | |||||
| Suffix Tree | Construct | |||||
| Skip List | Compound | |||||
| BST | Insert | |||||
| Remove | ||||||
| In-Order Traversal | ||||||
| Pre-Order Traversal | ||||||
| Post-Order Traversal | ||||||
| Depth | ||||||
| Compound | ||||||
| Heap | Compound | |||||
| Heapify | ||||||
| RB Tree | Construct | |||||
| Compound | ||||||
| B+ Tree | Compound | |||||
| Graph | Breadth-First Traversal | |||||
| Depth-First Traversal | ||||||
| DSU | Compound | |||||
| Bloom Filter | Compound | |||||
| DAWG | Compound |
| Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
| Array | Access | |||||
| Delete | ||||||
| Insert | ||||||
| Reverse | ||||||
| Search | ||||||
| Queue | Compound | |||||
| Stack | Compound | |||||
| LRU Cache | Cache | |||||
| Priority Queue | Compound | |||||
| Hashmap | Compound | |||||
| Trie | Compound | |||||
| Suffix Tree | Construct | |||||
| Skip List | Compound | |||||
| BST | Insert | |||||
| Remove | ||||||
| In-Order Traversal | ||||||
| Pre-Order Traversal | ||||||
| Post-Order Traversal | ||||||
| Depth | ||||||
| Compound | ||||||
| RB Tree | Construct | |||||
| Compound | ||||||
| B+ Tree | Compound | |||||
| Heap | Compound | |||||
| Heapify | ||||||
| Graph | Breadth-First Traversal | |||||
| Depth-First Traversal | ||||||
| DSU | Compound | |||||
| Bloom Filter | Compound | |||||
| DAWG | Compound |
| Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
| Array | Access | |||||
| Delete | ||||||
| Insert | ||||||
| Reverse | ||||||
| Search | ||||||
| Skip List | Compound | |||||
| Queue | Compound | |||||
| Stack | Compound | |||||
| LRU Cache | Compound | |||||
| Priority Queue | Compound | |||||
| Hashmap | Compound | |||||
| Trie | Compound | |||||
| Suffix Tree | Construct | |||||
| BST | Insert | |||||
| Remove | ||||||
| In-Order Traversal | ||||||
| Pre-Order Traversal | ||||||
| Post-Order Traversal | ||||||
| Depth | ||||||
| Compound | ||||||
| Heap | Compound | |||||
| Heapify | ||||||
| RB Tree | Construct | |||||
| Compound | ||||||
| B+ Tree | Compound | |||||
| Graph | Breadth-First Traversal | |||||
| Depth-First Traversal | ||||||
| DSU | Construct | |||||
| Bloom Filter | Compound | |||||
| DAWG | Compound |
A.6 DSR-Bench-spatial supplementary materials
Figure 3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn. These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section 4.2.
A.7 DSR-Bench-natural supplementary materials
Figure 4 illustrates our generation process for natural language prompts in DSR-Bench-natural. For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in A.7). This narrative is then paraphrased into five variants using GPT-4o. For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in A.7). In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
We use the same input distribution as in the DSR-Bench, replacing synthetic values (e.g. “enqueue 5”) with realistic values (e.g. “enqueue Yuki Lopez”). During prompt construction, we randomly sample one scenario description. For each data structure operation, we instantiate it with a synthetically generated value (e.g., a name) and sample one operation template to form a complete instance. All templates were reviewed by three human annotators to ensure clarity and unambiguous solvability.
Queue (children buying ice cream)
We construct a real-world scenario that implicitly models a queue: children lining up to buy ice cream from a truck. The enqueue operation corresponds to a child joining the end of the line, while the dequeue operation represents a child being served from the front. The scenario explicitly enforces the FIFO discipline by stating that no skipping is allowed.
BST (clinic appointments)
We construct a scenario in which a clinic uses a BST to store patient appointments, ordered by appointment time and tie-broken by the patient’s name. The insert operation adds a name, time) appointment to the tree, while the delete operation corresponds to a patient canceling their appointment. To retrieve all records, a pre-order traversal of the tree is performed.
Graph (galaxy traveling)
We create a scenario set in a galaxy where planets are connected by space tunnels. The task is to navigate a starship to visit as many planets as possible using depth-first search, starting from a given planet and visiting neighbors in lexicographical order.
A.8 Levenshtein distance: an auxiliary metric for DSR-Bench
In addition to the binary (0/1) accuracy reported in the main text, DSR,-Bench includes an optional evaluation metric based on Levenshtein distance, which measures the minimum number of single-character insertions, deletions, or substitutions needed to transform one string into another. As a continuous metric, it captures degrees of error that binary accuracy flattens. For instance, given the correct output [1,,3,,6], the prediction [3,,1,,6] is clearly closer than [0,,0,,0], and Levenshtein distance reflects that nuance.
However, this granularity can also blur important semantic distinctions. A syntactically well-formed but semantically incorrect output may still receive a high score, especially when the expected answer is long or formatted. When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein. For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table 25).
Output length further complicates cross-task comparison. Tasks with short, single-token outputs (e.g., Binary Search Tree depth) tend to show similar binary and Levenshtein scores (e.g., 0.66), while longer, multi-token outputs (e.g., Graph BFS) inflate Levenshtein scores (0.82) even when binary accuracy remains low (0.31). Relying on Levenshtein distance alone may thus give a misleading impression—for example, that the model performs well on BFS but poorly on tree depth—when binary accuracy indicates the opposite.
| Data structure | Operation | Binary | Levenshtein |
| Array | Access | ||
| Queue | Compound | ||
| Stack | Compound | ||
| LRU Cache | Cache | ||
| Priority Queue | Compound | ||
| Hashmap | Compound | ||
| Trie | Compound | ||
| Suffix Tree | Construct | ||
| Skip List | Compound | ||
| BST | Insert | ||
| Remove | |||
| Post-Order Traversal | |||
| Depth | |||
| Compound | |||
| Heap | Compound | ||
| Heapify | |||
| RB Tree | Construct | ||
| Compound | |||
| B+ Tree | Compound | ||
| Graph | Breadth-First Traversal | ||
| Depth-First Traversal | |||
| DSU | Compound | ||
| Bloom Filter | Compound | ||
| DAWG | Compound |
For these reasons, we report all results using binary accuracy and relegate Levenshtein evaluation to the toolkit. The implementation remains publicly available, as the metric can still offer a useful secondary perspective, particularly when comparing models on the same task and output length.

### v2

@@
-DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures
+Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures
 Abstract
-Large language models (LLMs) are increasingly deployed for real-world tasks that fundamentally involve data manipulation.
-A core requirement across these tasks is the ability to perform structural reasoning—that is, to understand and reason about data relationships.
-For example, customer requests require a temporal ordering, which can be represented by data structures such as queues.
-However, existing benchmarks primarily focus on high-level, application-driven evaluations without isolating this fundamental capability.
-To address this gap, we introduce DSR-Bench, a novel benchmark evaluating LLMs’ structural reasoning capabilities through data structures, which provide interpretable representations of data relationships.
-DSR-Bench includes 20 data structures, 35 operations, and 4,140 problem instances, organized hierarchically for fine-grained analysis of reasoning limitations.
-Our evaluation pipeline is fully automated and deterministic, eliminating subjective human or model-based judgments.
-Its synthetic nature also ensures scalability and minimizes data contamination risks.
-We benchmark nine state-of-the-art LLMs.
-Our analysis shows that instruction-tuned models struggle with basic multi-attribute and multi-hop reasoning.
-Furthermore, while reasoning-oriented models perform better, they remain fragile on complex and hybrid structures, with the best model achieving an average score of only 47% on the challenge subset.
-Crucially, models often perform poorly on multi-dimensional data and natural language task descriptions, highlighting a critical gap for real-world deployment.
+As large language models (LLMs) take on increasingly complex tasks, understanding their algorithmic reasoning abilities has become essential.
+However, existing evaluations focus on distinct and isolated tasks.
+We propose a unified diagnostic lens: structural reasoning—understanding and manipulating relationships like order, hierarchy, and connectivity.
+We introduce DSR-Bench, the first benchmark to systematically evaluate LLM structural reasoning through canonical data structures, which serve as interpretable, algorithmically meaningful abstractions.
+DSR-Bench spans 20 data structures, 35 operations, and 4,140 synthetically generated problem instances with minimal contamination.
+The benchmark’s hierarchical design pinpoints specific failure modes, while its fully automated evaluation ensures objective and consistent assessment.
+Benchmarking ten state-of-the-art LLMs reveals critical limitations: the top-performing model scores only 0.498 out of 1 on challenging instances.
+Three additional evaluation suites reveal further weaknesses: models perform poorly on spatial data and natural language scenarios, and fail to reason over their own generated code.
+DSR-Bench offers a principled diagnostic tool for structural reasoning, helping expose reasoning bottlenecks and guide the development of more capable and reliable LLMs.
 1 Introduction
-Large language models (LLMs) have demonstrated impressive capabilities across practical tasks such as database management [62], coding [3, 58], and scheduling [20].
-These tasks differ in domain, but they share a core requirement: the ability to understand and manipulate data according to relationships such as order, hierarchy, and connectivity—a capability we refer to as structural reasoning.
-Structural reasoning is an essential, yet often implicit, requirement in problem solving.
-For instance, when using an LLM to plan a trip, the system may need to interpret a map of cities as a graph, construct an itinerary by finding a path through that graph, and prioritize visits using a queue.
-As LLMs are increasingly deployed in real-world applications, particularly safety-critical ones such as robotics [39], healthcare [38], and autonomous systems [14, 1], it becomes imperative to rigorously evaluate their structural reasoning capabilities to identify potential failure modes prior to deployment.
-Furthermore, LLMs are known to struggle with complex forms of reasoning, such as spatial reasoning [22, 55], underscoring the timely need for a systematic evaluation of their structural reasoning abilities.
-However, existing benchmarks offer limited insight into structural reasoning.
-Most evaluations focus on high-level, domain-specific tasks such as mathematical problem solving [28, 25, 50], competitive programming [16], or STEM-related questions [15].
-These tasks are typically multi-step and complex, making it difficult to isolate structural reasoning performance from other reasoning abilities.
-As a result, a critical gap remains: there are no targeted, fine-grained benchmarks for evaluating structural reasoning, which is a core skill underlying many downstream applications
-To address this gap, we propose evaluating structural reasoning through the lens of data structures.
-Data structures offer a controlled and interpretable framework for modeling diverse data relationships: arrays for sequences, stacks and queues for temporal logic, trees for hierarchies, and graphs for complex networks.
-Since structural reasoning manifests as operations on these canonical structures, data structure tasks provide a principled way to probe LLM reasoning abilities at a granular level.
-Such a fine-grained evaluation can not only identify reliable applications of LLMs for safe deployment, but also pinpoint areas where reasoning breaks down and how it might be improved.
-Building on this insight, we introduce DSR-Bench, a benchmark designed to isolate and evaluate LLM structural reasoning via data structure manipulation tasks.
-DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories, see Figure 1) and 35 operations across three length types (short, medium, and long).
-We also introduce DSR-Bench-challenge, a subset of DSR-Bench consisting of data structures with particularly complex designs.
-Each task probes whether the model can construct, maintain, and reason about data relationships within an interpretable and automatically verifiable setting.
-DSR-Bench offers several key advantages for evaluating structural reasoning:
-Hierarchical organization.
-Tasks are organized by increasing structural complexity, enabling a fine-grained analysis of specific reasoning skills.
-Within each category, we design a range of tasks to isolate different sources of complexity, allowing structural reasoning to be broken down into progressively more challenging tasks.
-This approach precisely identifies the specific types of data relationships that pose difficulties for LLMs.
-Deterministic evaluation.
-Each data structure task has a concise and well-defined final state as its ground-truth label, allowing for deterministic and unambiguous scoring.
-Unlike open-ended tasks, this design supports fully automated evaluation without the need for human or model-based judgment, resulting in a fairer and more objective evaluation pipeline [4, 10, 54].
-We benchmark nine state-of-the-art LLMs, including both instruction-tuned and reasoning models. Our analysis reveals the following key insights:
-1.
-Instruction-tuned models struggle with multi-attribute and multi-hop reasoning.
-Their accuracy drops sharply on tasks involving multi-attribute data (such as hashmaps) and tasks requiring the maintenance of multi-hop properties (such as red-black trees).
-Prompting strategies that encourage intermediate steps can improve performance, but techniques like CoT tend to be most effective when the data structure is less standard.
-2.
-Reasoning models still struggle with complex structures and non-standard tasks.
-Despite outperforming instruction-tuned models, reasoning models perform poorly on structurally complex tasks, with the best scoring only 47% on DSR-Bench-challenge.
-Some also fail when task instructions conflict with what are likely learned assumptions (e.g., consistently predicting tree depth as the number of edges instead of nodes, despite explicit instructions).
-These errors suggest a reliance on entrenched priors, limiting the model’s ability to adapt to user-defined constraints.
-3.
-Model performance declines significantly on complex spatial data structures.
-Accuracy drops as data dimensionality increases and degrades further when data is non-uniform.
-These patterns suggest a reliance on memorization over robust generalization or genuine reasoning.
-4.
-LLMs perform worse when data structure tasks are presented in natural language.
-Translating formal task descriptions into natural language leads to a marked drop in performance, posing a barrier to the practical deployment of LLMs for real-world problem solving.
-We release all code at https://github.com/dransyhe/DSR-Bench and all data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 to support full reproducibility.
-We welcome community engagement and collaboration.
+As large language models (LLMs) tackle increasingly complex real-world challenges, strengthening their algorithmic reasoning abilities is essential for interpretability, safety, and efficiency (Eberle et al., 2025).
+Yet scaling alone is reaching its limits: models face data scarcity and diminishing returns (Villalobos et al., 2024).
+A complementary path is to study the algorithms LLMs learn and employ, which can reveal their internal representations and inspire algorithm-centric architectures that embed reasoning capabilities directly into model design (Bounsi et al., 2024; Eberle et al., 2025).
+The call to prioritize algorithmic reasoning has intensified as the field shifts toward evaluating general reasoning without external tools.
+While LLMs can solve many tasks through code generation and execution, recent initiatives like Gemini-Deep-Think’s and OpenAI’s IMO competitions (Luong & Lockhart, 2025; Wei, 2025) explicitly prohibit coding and proof assistance.
+These efforts highlight a growing emphasis on end-to-end reasoning as a critical step toward artificial general intelligence.
+Despite rapid progress, existing evaluations of algorithmic reasoning remain fragmented, each probing only a single facet, such as arithmetic operations (Zhou et al., 2022; 2024; Lee et al., 2024), simulation of textbook algorithms (Markeeva et al., 2024), or domain-specific tasks like graph problems (Wang et al., 2023a; Fatemi et al., 2024).
+Meanwhile, interpretability research, such as mechanistic interpretability (Olah et al., 2020; Wang et al., 2023b), examines internal model representations rather than reasoning over algorithmic structures.
+What is missing is a unified framework that bridges these threads, offering interpretable diagnostics of algorithmic reasoning across diverse problem types.
+To address this gap, we propose to evaluate algorithmic reasoning through the lens of data structures—fundamental abstractions that organize information and govern algorithmic operations.
+Data structures provide an ideal evaluation framework because they are both systematic (covering diverse relationship types) and interpretable (offering clear diagnostic insights).
+They span the spectrum of relationships underlying algorithmic thinking: arrays represent sequences, stacks and queues capture temporal ordering, trees encode hierarchies, and graphs represent complex networks.
+This diversity enables us to holistically assess whether models can construct, maintain, and manipulate different types of relationships—the core capability we term structural reasoning.
+Structural reasoning underlies many real-world applications.
+Trip planning requires interpreting maps as graphs and managing priorities with queues, while supply chain optimization depends on hierarchical resource allocation and temporal sequencing.
+In safety-critical domains such as robotics (Sado et al., 2023) and healthcare (Sadeghi et al., 2024), failures in structural reasoning abilities can have serious consequences, making systematic evaluation crucial before deployment.
+Contributions
+We introduce DSR-Bench, the first benchmark to systematically isolate and evaluate LLM structural reasoning through data structure manipulation tasks.
+DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories; see Figure˜1) and 35 operations across three difficulty levels (short, medium, and long) and five evaluation suites (main, challenge, spatial, natural, and code).
+Each task provides interpretable, automatically verifiable assessments of whether models can construct, maintain, and reason about relationships.
+The highlights and main takeaways of DSR-Bench are:
+Hierarchical and diagnostic design. Simpler tasks (e.g., queue operations) serve as prerequisites for complex ones (e.g., graph traversal). This enables fine-grained localization of failure modes.
+Deterministic, contamination-resistant evaluation. Synthetic data minimizes contamination risk. The evaluation is deterministic and fully automated, supporting objective and fair assessment.
+Five evaluation suites covering diverse reasoning settings.
+main: Canonical data structure tasks reveal core reasoning gaps—e.g., instruction-tuned models struggle with multi-attribute and multi-hop reasoning, while reasoning models fail to override memory for non-standard tasks with user-defined constraints.
+Simpler prompts consistently help, but CoT requires careful design.
+challenge: Hybrid and compositional structures remain difficult, with the best model scoring only 0.498 out of 1, revealing reasoning limitations of frontier models.
+spatial: On multi-dimensional data, the common format found in real-world applications, models struggle as dimensionality increases or when data follows non-uniform distributions.
+natural: In realistic natural language scenarios, models struggle to extract structure and navigate ambiguity, exposing a gap to real-world deployment.
+code: An auxiliary probe contrasting reasoning with and without tool use.
+Models rarely benefit from reasoning over their own generated code; external interpreters help on familiar tasks but consistently fail on non-standard or natural-language variants.
+Comprehensive empirical analysis.
+We evaluate ten state-of-the-art LLMs and analyze prompting strategies, distribution shifts, and qualitative error patterns (e.g., implicit priors, instruction-following failures).
+These findings provide actionable insights for developing architectures and training methods that better capture algorithmic reasoning.
+To welcome community engagement and collaboration, we release all code at https://github.com/dransyhe/DSR-Bench and data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 for full reproducibility.
+2 Related work
+Algorithmic reasoning with LLMs
+Prior works target distinct but isolated facets of algorithmic reasoning, often limited in scope: arithmetic tasks and length generalization in transformers (Zhou et al., 2022; 2024; Lee et al., 2024), or specific graph problems like cycle detection and connectivity (Wang et al., 2023a; Fatemi et al., 2024).
+CLRS-Text (Markeeva et al., 2024) evaluates whether LLMs can simulate 30 classical algorithms (Veličković et al., 2022; Cormen et al., 2009).
+We take a holistic and foundational approach using data structures, which serve as interpretable building blocks of algorithms and support fine-grained diagnosis of specific reasoning failures to inform future research.
+Reasoning benchmarks
+Existing LLM reasoning benchmarks are predominantly high-level and domain-specific, targeting math (Cobbe et al., 2021; Liu et al., 2024a; b), STEM (Hendrycks et al., 2021), and logic puzzles (White et al., 2025; Giadikiaroglou et al., 2024).
+These often require complex responses with intertwined reasoning steps, relying on subjective human or LLM-based evaluation (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024).
+We focus on structural reasoning, an implicit requirement underlying problem-solving across domains.
+By using data structures as clear abstractions of these relationships, we isolate structural reasoning from domain-specific complexities.
+Coding benchmarks
+Coding benchmarks evaluate how well LLMs write syntactically correct code or function as coding agents (Chen et al., 2021; Zheng et al., 2023; Jimenez et al., 2024; Jain et al., 2025; White et al., 2025; Aider-AI, 2025), typically requiring external interpreters for verification.
+While useful, these benchmarks conflate reasoning with tool execution and are limited to domains where coding applies.
+In contrast, we target general reasoning independent of external tools, reflecting the broader goal of assessing progress toward general intelligence.
+To bridge these perspectives, we include the code suite to probe whether code generation aids structural reasoning.
 3 DSR-Bench: the Data Structure Reasoning Benchmark
 In DSR-Bench, we propose a taxonomy of data structures grouped into six categories based on the types of data relationships they encode: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid.
-This taxonomy captures the diversity of relational patterns found in real-world data and enables systematic evaluation of LLM capabilities across distinct relationship types.
-Within each category, we include a spectrum of data structures that vary in complexity and operational constraints, allowing us to isolate specific reasoning challenges and precisely diagnose where models struggle.
-2.1 Data structure tasks
-DSR-Bench includes the following data structures and tasks. A detailed description of each data structure is provided in Section A.2.
+This taxonomy captures the diversity of relational patterns found in real-world data, enabling systematic evaluation across distinct reasoning challenges.
+3.1 Tasks
+DSR-Bench includes the following data structures and tasks, further detailed in Section˜A.2.
 Linear (Sequential): This category includes array and its operations: access, insert, delete, reverse, and search.
-Linear structures are fundamental as they introduce ordered relationships, enabling reasoning about position, sequence, and iteration, and they serve as a foundation for understanding more complex data abstractions.
+Linear structures introduce ordered relationships, enabling reasoning about position, sequence, and iteration.
+They serve as a foundation for more complex data abstractions.
 Temporal (Time-based ordering): Temporal structures include stack, queue, priority queue, and skip list, which operate under last-in-first-out, first-in-first-out, priority-based, or probabilistic rules.
 Their operations are compound actions of insertions and deletions.
-Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers, and introduces an additional ordering constraint beyond simple linear sequencing.
+Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers.
 Associative (Key-value mapping): We evaluate hashmap, trie, and suffix tree through tasks focused on construction and compound insertion-deletion operations.
-These tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures, which are core skills for efficient lookup and structured access in systems such as databases.
+Tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures.
+These skills are core for efficient lookup and structured access in systems such as databases.
 Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, with tasks involving traversals and compound operations.
 These structures are common in file systems and database indexing.
-These tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
-Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU).
-Graph tasks focus on breadth- and depth-first traversal, testing a model’s ability to reason about connectivity and explore many-to-many relationships, such as user interactions in social networks.
-DSU tasks centre on union-find operations, which are fundamental for dynamically managing group membership and identifying connected components in partitioned data.
-Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple relational principles.
-This category includes LRU cache, which integrates temporal and memory constraints; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph.
-Tasks involve compound insertions and deletions, testing whether models can reason over combined constraints and generalize beyond individual structural concepts.
-2.2 Data generation
-We design a diverse set of operation tasks for each data structure, as detailed in Table 8.
-Tasks fall into three categories: construction, inspection (e.g., access, pre/in/post-order traversal), and manipulation (e.g., insert, delete).
-In addition to these atomic operations, which serve as fundamental building blocks and allow fine-grained diagnosis of specific failure modes, DSR-Bench also includes compound operations.
-These are composed of a sequence of manipulations (e.g., insert and delete) and are designed to test a model’s ability to maintain structural integrity across multiple steps.
-All data is synthetically generated, which is efficient and less susceptible to training data contamination [56, 51].
-Numerical inputs are sampled uniformly from the range 0 to 100, while string inputs are composed of lowercase English letters.
-For each task, ground-truth outputs are generated by implementing the corresponding data structure and executing the prescribed operations.
-The expected format of the final state is explicitly specified in the prompt.
-For simple structures like arrays, the final state is a sequence of elements.
-For more complex structures such as trees and graphs, the final state is represented using a canonical traversal, such as pre-order or breadth-first.
-The modular design of our benchmark also supports easy incorporation of new data structure types.
-2.3 Suites
-We categorize each task into three levels based on input length: short (5–10), medium (11–20), and long (21–30), to evaluate length generalization.
-For atomic operations, length corresponds to the number of input elements, such as list items or tree nodes.
-For compound operations, tasks start from an empty structure, and length is defined by the number of sequential operations applied.
-We also curated a DSR-Bench-challenge subset highlighting data structures with particularly complex designs, aimed at pushing the limits of state-of-the-art models.
-To further assess structural reasoning capabilities, we also introduce two extension suites: DSR-Bench-spatial, focusing on spatial reasoning on multi-dimensional inputs (details see Section 4.2), and DSR-Bench-natural, evaluating reasoning over natural language descriptions (details see Section 4.3).
+Tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
+Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU) tasks.
+Graph tasks cover breadth-first and depth-first traversal, testing the ability to reason about connectivity and explore many-to-many relationships, as commonly found in social networks.
+DSU tasks evaluate union–find operations, which are essential for network connectivity analysis and clustering algorithms.
+Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple forms of structural reasoning.
+This category includes LRU cache, which integrates temporal and memory management; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph.
+Tasks test whether models can compose different principles and generalize beyond individual structures.
+Hierarchical organization
+Tasks are organized by increasing complexity for fine-grained diagnosis of reasoning failures.
+For example, success on list but failure on hashmap isolates the issue to multi-attribute access rather than linear sequencing.
+Similarly, since tree traversal depends on queue or stack operations, testing these primitives becomes a prerequisite.
+Operations
+We design a diverse set of operation tasks for each data structure (full list in Table˜10), spanning three categories: construction, inspection (e.g., access, traversal), and manipulation (e.g., insert, delete).
+Beyond these atomic operations, DSR-Bench includes compound operations: sequences of manipulations (e.g., insert, insert, delete, …) designed to test whether models can maintain structural integrity across multiple steps.
+Difficulty levels
+We assign tasks to three difficulty levels based on input length: short (5–10), medium (11–20), and long (21–30), to assess length generalization.
+For atomic operations, length refers to the number of input elements (e.g., list items, tree nodes).
+For compound operations, tasks begin with an empty structure, and length is defined by the number of sequential operations.
+Evaluation suites
+Beyond the main suite, we curated a challenge subset targeting particularly complex structures to stress-test advanced reasoning capabilities.
+To more comprehensively evaluate structural reasoning beyond canonical data structures, we introduce three additional suites: spatial, which evaluates reasoning on multi-dimensional data; natural, which tests reasoning when tasks are embedded in realistic natural language scenarios; and code, which assesses whether models can leverage code generation for structural reasoning.
+Each suite examines how well structural reasoning capabilities transfer to scenarios critical for real-world deployment.
 3.2 Prompt design
@@
 Each prompt follows a consistent format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final computed outcome.
-Following recommended practices in prompt engineering [35], we also append the instruction “Answer the question in <token> tokens” to guide models toward producing concise outputs within a specified token budget.
-We also implement five different prompting methods for each task, as we detail in Section 3.2.
-2.5 Structured output
-A key advantage of our tasks is that model outputs can be deterministically verified using automated evaluation.
-This is possible because the final states of data structures can be represented as sequences of numbers or strings, enabling precise comparison.
-Unlike benchmarks involving proof writing or complex math problems, which often yield long and ambiguous outputs requiring subjective human or LLM-based judging [4, 10, 54], our evaluation framework is more objective and reproducible.
-For automatic verification, models must produce outputs matching predefined JSON schemas enforced with Pydantic.
-For instance, a Binary Search Tree (BST) pre-order traversal returns a list[int].
-We ensure that output formats are easy to parse: about 70% of the tasks yield a 1D integer array.
-2.6 Scoring system
-We evaluate each model by comparing its output to the ground-truth answer.
-Each correct output scores 1, otherwise 0.
-The overall performance is reported as the average accuracy across all categories and length levels.
-To ensure deterministic outputs and eliminate ambiguity during evaluation, every task admits a single, well-defined solution.
-Potential sources of nondeterminism, such as definitions of hash function and tie-breaking rules, are explicitly specified in the prompt and controlled during problem generation.
-Additionally, we run extensive tests across all tasks and manually inspect failure cases to verify that errors are not caused by ambiguity in the question statements or task specifications.
-3 Experiment settings
-In this section, we describe our experimental settings. Section 3.1 outlines the models evaluated, and Section 3.2 details the prompting methods used.
-3.1 Models
-We evaluate a broad set of LLMs across two categories: general-purpose instruction-tuned models and models designed for advanced reasoning.
-From each major provider, we select the latest publicly available flagship models.
-Each prompt is evaluated three times, and we report both the average and standard deviation, resulting in a total of 166,230 query evaluations.
-Instruction-tuned LLMs are widely used in real-world applications due to their low latency, reduced cost, and scalability.
-Although not explicitly trained for algorithmic or mathematical reasoning, their ability to process structured data is critical in practical settings.
-We test GPT-4.1-2025-04-14 (OpenAI) [36], Llama3.3-70B (Meta) [13], Gemini-2.0-Flash-001 (Google DeepMind) [45] , Claude-3-5-Sonnet-20241022 (Anthropic) [2], and DeepSeek-V3 (DeepSeek) [7].
-We also evaluate a suite of models optimized for advanced reasoning.
-These models are explicitly trained to handle complex, multi-step tasks by maintaining intermediate states, including: o4-mini-2025-04-16 (OpenAI) [34], Gemini-2.5-Pro-Preview-03-25 (Google DeepMind) [45], Claude-3-7-Sonnet-20250219 (Anthropic) [2], and DeepSeek-R1 (DeepSeek) [7].
-3.2 Prompts
-To examine how prompting strategies affect model performance on data structure tasks, we evaluate five configurations.
-Unlike reasoning models with internal mechanisms for multi-step inference (e.g., explicit reasoning tokens), instruction-tuned models are especially sensitive to prompt formulation.
-Our prompting configurations include: (i) Stepwise, which adds a “steps” field to the JSON output; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single reasoning example with a final output; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text.
-See Section A.3 for examples.
+Following recommended practices in prompt engineering (OpenAI, 2025), we also append the instruction “Answer the question in <number> tokens” to guide models toward producing concise outputs within a specified token budget.
+We also implement five different prompting strategies for each task, as we detail in Section˜A.3.
+3.3 Pipeline
+Data generation
+All data is synthetically generated to ensure efficiency, scalability, and minimal risk of training data contamination (Zhang et al., 2024a; Xu et al., 2024a).
+Each data structure and its operations are programmatically implemented to produce ground-truth outputs.
+Inputs are randomly sampled—numbers uniformly from 0–100 and strings from lowercase English letters.
+This modular approach enables accurate, reproducible evaluation and easy integration of new data structure types.
+Automated evaluation
+The evaluation pipeline is fully objective and reproducible through automated and deterministic verification.
+This avoids subjective judging in proof-writing or math benchmarks (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024).
+We integrate Structured Output using a predefined JSON Schema (natively supported by most models or via the Instructor library) and Pydantic, ensuring strict schema adherence.
+For example, a BST pre-order traversal must return a list[int].
+Across 1,624 trials with nine models and six structures, we observed zero schema violations (Table˜33 in Appendix), confirming robustness of Structured Output.
+Scoring system
+We compare model outputs to ground-truth answers, scoring 1 for correct and 0 for incorrect responses.
+Overall performance is reported as average accuracy across categories and difficulty levels from three runs.
+To ensure determinism, each task has a single well-defined solution.
+Any potential ambiguity (hash functions, tie-breaking rules) is explicitly specified in the prompts.
+We include an ablation with Levenshtein distance as a partial scoring metric in Section˜A.10.
 4 Evaluation
-In this section, we present experimental results evaluating all nine LLMs and highlight their strengths and weaknesses across structural reasoning tasks on DSR-Bench and DSR-Bench-challenge (Section 4.1), DSR-Bench-spatial (Section 4.2), and DSR-Bench-natural (Section 4.3).
+Models
+We evaluate ten state-of-the-art LLMs across two categories: instruction-tuned models and reasoning models.
+We select the latest flagship models from each major provider, with each problem evaluated three times (166,230 total evaluations).
+Instruction-tuned models are widely deployed due to their efficiency and scalability, making their structural reasoning capabilities practically important: we cover GPT-4.1-2025-04-14 (OpenAI et al., 2024), Llama3.3-70B (Grattafiori et al., 2024), Gemini-2.0-Flash-001 (Team et al., 2024), Claude-3-5-Sonnet-20241022 (Anthropic, 2024), and DeepSeek-V3 (DeepSeek-AI et al., 2025).
+Reasoning models are explicitly trained for complex, multi-step reasoning: we evaluate GPT-5-2025-08-07 with medium thinking effort (OpenAI, 2025), o4-mini-2025-04-16 (OpenAI, 2024), Gemini-2.5-Pro (stable) (Team et al., 2024), Claude-3-7-Sonnet-20250219 (Anthropic, 2024), and DeepSeek-R1 (DeepSeek-AI et al., 2025).
+Prompting strategies
+We also study the impact of prompting strategies on data structure tasks.
+Unlike reasoning models with internal multi-step inference (e.g., reasoning tokens), instruction-tuned models are particularly sensitive to prompt formulation.
+We evaluate five prompting strategies: (i) Stepwise, which adds a “steps” field to the output JSON schema; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single example with intermediate reasoning steps; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text.
+See Section˜A.3 for examples.
 4.1 Can LLMs understand and manipulate data structures?
-In this section, we present our experimental results of LLMs on DSR-Bench (including DSR-Bench-challenge).
-In Section 4.1.1, we discuss the results of the five instruction-tuned models, and in Section 4.1.2, we highlight key observations on the four reasoning models.
-| Relationship | Data Structure | o4- mini | Gemini- 2.5-Pro | Claude-3. 7-Sonnet | DeepSe ek-R1 | GPT- 4.1 | Gemini- 2.0-Flash | Claude-3. 5-Sonnet | DeepSe ek-V3 | Llama 3.3-70B |
-| Linear | Array | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 |
-| Category avg. | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 | |
-| Temporal | Stack | 1.00 | 1.00 | 0.97 | 0.99 | 0.55 | 0.36 | 0.99 | 0.41 | 0.04 |
-| Queue | 1.00 | 1.00 | 1.00 | 0.98 | 0.55 | 0.36 | 0.99 | 0.43 | 0.25 | |
-| LRU | 1.00 | 1.00 | 0.30 | 0.16 | 0.85 | 0.78 | 0.82 | 0.01 | 0.50 | |
-| Priority Queue | 0.55 | 0.51 | 0.28 | 0.65 | 0.25 | 0.16 | 0.33 | 0.20 | 0.08 | |
-| Category avg. | 0.89 | 0.84 | 0.64 | 0.69 | 0.55 | 0.42 | 0.79 | 0.26 | 0.22 | |
-| Associative | Hashmap | 0.51 | 0.64 | 0.63 | 0.33 | 0.06 | 0.10 | 0.16 | 0.00 | 0.00 |
-| Trie | 0.68 | 0.80 | 0.08 | 0.49 | 0.18 | 0.17 | 0.49 | 0.02 | 0.00 | |
-| Suffix Tree | 0.73 | 0.85 | 0.91 | 0.96 | 0.00 | 0.01 | 0.08 | 0.67 | 0.00 | |
-| Skip List | 0.62 | 0.73 | 0.75 | 0.69 | 0.07 | 0.06 | 0.42 | 0.02 | 0.01 | |
-| Category avg. | 0.63 | 0.76 | 0.59 | 0.62 | 0.08 | 0.09 | 0.29 | 0.18 | 0.00 | |
-| Hierarchical | BST | 0.86 | 0.86 | 0.64 | 0.73 | 0.59 | 0.43 | 0.71 | 0.58 | 0.34 |
-| Heap | 0.68 | 0.61 | 0.40 | 0.48 | 0.20 | 0.10 | 0.27 | 0.15 | 0.07 | |
-| RB tree | 0.65 | 0.77 | 0.30 | 0.62 | 0.12 | 0.12 | 0.46 | 0.09 | 0.05 | |
-| B+ tree | 0.97 | 0.95 | 0.38 | 0.31 | 0.23 | 0.12 | 0.23 | 0.08 | 0.01 | |
-| K-D Tree | 0.47 | 0.71 | 0.00 | 0.45 | 0.00 | 0.01 | 0.03 | 0.00 | 0.00 | |
-| K-D Heap | 0.10 | 0.05 | 0.05 | 0.11 | 0.04 | 0.04 | 0.05 | 0.03 | 0.01 | |
-| Category avg. | 0.64 | 0.78 | 0.33 | 0.45 | 0.20 | 0.14 | 0.34 | 0.16 | 0.08 | |
-| Network | Graph | 0.87 | 0.94 | 0.11 | 0.67 | 0.15 | 0.05 | 0.06 | 0.06 | 0.02 |
-| DSU | 0.98 | 1.00 | 0.99 | 1.00 | 0.02 | 0.00 | 0.02 | 0.82 | 0.00 | |
-| Geom Graph | 0.16 | 0.13 | 0.02 | 0.13 | 0.07 | 0.02 | 0.02 | 0.01 | 0.02 | |
-| Category avg. | 0.67 | 0.69 | 0.38 | 0.60 | 0.08 | 0.03 | 0.03 | 0.30 | 0.01 | |
-| Hybrid | Bloom Filter | 0.61 | 0.89 | 0.16 | 0.74 | 0.04 | 0.04 | 0.04 | 0.04 | 0.02 |
-| DAWG | 0.25 | 0.07 | 0.06 | 0.14 | 0.05 | 0.06 | 0.07 | 0.06 | 0.01 | |
-| Category avg. | 0.43 | 0.48 | 0.11 | 0.44 | 0.05 | 0.05 | 0.06 | 0.05 | 0.02 | |
-| Score | Overall avg. | 0.72 | 0.76 | 0.51 | 0.63 | 0.31 | 0.27 | 0.41 | 0.32 | 0.17 |
+In this section, we present our main experimental results on DSR-Bench, including main and its challenging subset challenge.
+We discuss insights from five instruction-tuned models in Section˜4.1.1 and from five reasoning models in Section˜4.1.2.
+| Relationship | Data Structure | GPT-5 (med) | o4- mini | Gemini- 2.5-Pro | Claude-3. 7-Sonnet | DeepSe ek-R1 | GPT- 4.1 | Gemini- 2.0-Flash | Claude-3. 5-Sonnet | DeepSe ek-V3 | Llama 3.3-70B |
+| Linear | Array | 1.00 | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 |
+| Category avg. | 1.00 | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 | |
+| Temporal | Stack | 1.00 | 1.00 | 1.00 | 0.97 | 0.99 | 0.55 | 0.36 | 0.99 | 0.41 | 0.04 |
+| Queue | 1.00 | 1.00 | 1.00 | 1.00 | 0.98 | 0.55 | 0.36 | 0.99 | 0.43 | 0.25 | |
+| LRU | 1.00 | 1.00 | 1.00 | 0.30 | 0.16 | 0.85 | 0.78 | 0.82 | 0.01 | 0.50 | |
+| Priority Queue | 0.52 | 0.55 | 0.51 | 0.28 | 0.65 | 0.25 | 0.16 | 0.33 | 0.20 | 0.08 | |
+| Category avg. | 0.88 | 0.89 | 0.84 | 0.64 | 0.69 | 0.55 | 0.42 | 0.79 | 0.26 | 0.22 | |
+| Associative | Hashmap | 0.87 | 0.51 | 0.28 | 0.63 | 0.33 | 0.06 | 0.10 | 0.16 | 0.00 | 0.00 |
+| Trie | 0.94 | 0.68 | 0.62 | 0.08 | 0.49 | 0.18 | 0.17 | 0.49 | 0.02 | 0.00 | |
+| Suffix Tree | 0.98 | 0.73 | 0.90 | 0.91 | 0.96 | 0.00 | 0.01 | 0.08 | 0.67 | 0.00 | |
+| Skip List | 0.68 | 0.62 | 0.78 | 0.75 | 0.69 | 0.07 | 0.06 | 0.42 | 0.02 | 0.01 | |
+| Category avg. | 0.87 | 0.63 | 0.65 | 0.59 | 0.62 | 0.08 | 0.09 | 0.29 | 0.18 | 0.00 | |
+| Hierarchical | BST | 1.00 | 0.86 | 0.97 | 0.64 | 0.73 | 0.59 | 0.43 | 0.71 | 0.58 | 0.34 |
+| Heap | 0.61 | 0.68 | 0.38 | 0.40 | 0.48 | 0.20 | 0.10 | 0.27 | 0.15 | 0.07 | |
+| RB tree | 0.76 | 0.65 | 0.49 | 0.30 | 0.62 | 0.12 | 0.12 | 0.46 | 0.09 | 0.05 | |
+| B+ tree | 0.98 | 0.97 | 0.97 | 0.38 | 0.31 | 0.23 | 0.12 | 0.23 | 0.08 | 0.01 | |
+| K-D Tree | 0.85 | 0.47 | 0.59 | 0.00 | 0.45 | 0.00 | 0.01 | 0.03 | 0.00 | 0.00 | |
+| K-D Heap | 0.10 | 0.10 | 0.10 | 0.05 | 0.11 | 0.04 | 0.04 | 0.05 | 0.03 | 0.01 | |
+| Category avg. | 0.72 | 0.64 | 0.58 | 0.33 | 0.45 | 0.20 | 0.14 | 0.34 | 0.16 | 0.08 | |
+| Network | Graph | 0.96 | 0.87 | 0.78 | 0.11 | 0.67 | 0.15 | 0.05 | 0.06 | 0.06 | 0.02 |
+| DSU | 1.00 | 0.98 | 0.97 | 0.99 | 1.00 | 0.02 | 0.00 | 0.02 | 0.82 | 0.00 | |
+| Geom Graph | 0.16 | 0.16 | 0.13 | 0.02 | 0.13 | 0.07 | 0.02 | 0.02 | 0.01 | 0.02 | |
+| Category avg. | 0.71 | 0.67 | 0.63 | 0.38 | 0.60 | 0.08 | 0.03 | 0.03 | 0.30 | 0.01 | |
+| Hybrid | Bloom Filter | 0.77 | 0.61 | 0.86 | 0.16 | 0.74 | 0.04 | 0.04 | 0.04 | 0.04 | 0.02 |
+| DAWG | 0.34 | 0.25 | 0.23 | 0.06 | 0.14 | 0.05 | 0.06 | 0.07 | 0.06 | 0.01 | |
+| Category avg. | 0.56 | 0.43 | 0.55 | 0.11 | 0.44 | 0.05 | 0.05 | 0.06 | 0.05 | 0.02 | |
+| Score | Overall avg. | 0.79 | 0.72 | 0.71 | 0.51 | 0.63 | 0.31 | 0.27 | 0.41 | 0.32 | 0.17 |
 4.1.1 Instruction-tuned models
 Instruction-tuned models struggle with multi-attribute reasoning.
-As shown in Table 1, these models show sharp accuracy drops in tasks involve elements with multiple attributes.
+As shown in Table˜1, these models show sharp accuracy drops in tasks involving elements with multiple attributes.
 For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority.
@@
 Multi-hop reasoning in hierarchical or network structures remains a key challenge.
-We see from Table 1 that, while models perform reliably on binary search trees, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across multiple ancestral levels.
-Performance declines further on B+ trees, which require reasoning over wider spans of child pointers, and on graph traversal tasks involving many-to-many relationships.
-These results underscore a persistent limitation: instruction-tuned LLMs struggle to maintain and manipulate context beyond local neighborhoods.
+We see from Table˜1 that, while models perform reliably on BST, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across ancestral levels.
+Performance declines further on B+ trees, which requires reasoning over wider spans of child pointers, and on graph traversal tasks with many-to-many relationships.
+Manual inspection of reasoning traces reveals failures to retain earlier information: in Geometric Graph-long, all GPT-4.1 errors stemmed from dropping nodes during intermediate steps.
 | Model | Stepwise | 0-CoT | CoT | 3-shot | None |
@@
 Prompting can help, but only when carefully designed.
-As shown in Table 2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial.
-Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures.
-A representative case is suffix-tree construction: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section A.5).
+As shown in Table˜2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial.
+Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section˜A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures, but require careful design.
+A representative case is suffix-tree: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section˜A.5).
+We include more analysis of CoT to provide meaningful takeaways for practitioners in Section˜A.6.
 4.1.2 Reasoning models
 | Model | Score | Priority Queue | Suffix Tree | Trie | Skip List | Heap | Red-bla ck tree | B+ Tree | K-D Tree | K-D Heap | DSU | Geom Graph | Bloom Filter | DAWG |
+| GPT-5 | 0.50 | 0.52 | 0.98 | 0.92 | 0.51 | 0.71 | 0.585 | 0.98 | 0.85 | 0.10 | 1.00 | 0.19 | 0.47 | 0.00 |
 | Gemini-2.5-Pro | 0.47 | 0.24 | 0.89 | 0.59 | 0.61 | 0.62 | 0.60 | 0.96 | 0.67 | 0.00 | 0.99 | 0.00 | 0.69 | 0.00 |
@@
 Reasoning models remain brittle on complex and spatial data structures.
-As shown in Table 1, reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures.
-However, overall score remains below 0.5 on DSR-Bench-challenge (Table 3), in particular for long tasks and complex data structures.
-For example, the highest score on heap is only 0.62, despite it being a widely used structure in scheduling and resource allocation.
-Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Appendix A.4).
-To further probe these limitations, we introduce DSR-Bench-spatial, described in Section 4.2.
+From Table˜1, we see reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures.
+However, the overall score remains below 0.5 on challenge in Table˜3, in particular for long tasks and complex data structures.
+For example, the highest score on skip list is only 0.61, despite its prevalence in introductory-level textbooks and its wide use in dictionaries and maps.
+Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Section˜A.4).
+To further probe these limitations, we introduce spatial, described in Section˜4.2.
 Implicit priors may hinder instruction following.
-We observe that while GPT-4.1 scores 0.66 on the binary search tree depth task (short), o4-mini performs much worse, with a score of just 0.01 (Section A.4).
-Manual inspection of incorrect outputs reveals that o4-mini consistently interprets depth as the number of edges rather than nodes, despite explicit instructions.
-Similarly, in an ablation on K-D Heap (Table 5), switching from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons.
-Directly quering o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys.
-These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors, limiting their reliability on tasks with strict user-defined constraints.
+In an ablation on K-D Heap in Table˜5, switching the tie-breaking rule from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons.
+Directly querying o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys.
+These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors learned from training, limiting their reliability on tasks with user-defined constraints.
+4.2 Can LLMs reason structurally on spatial data?
+Real-world data is often represented in high-dimensional feature spaces.
+To assess whether LLMs can reason over such spatial data, we extend the benchmark with the spatial suite, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space.
+These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics.
+Given the complexity of these tasks, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps.
 | K-D Heap | K-D Tree | Geom. Graph | ||||
@@
 | 5 | 0.90 | 0.21 |
-4.2 Can LLMs reason structurally on spatial data?
-Real-world data is often represented in multi-dimensional feature spaces.
-To assess whether LLMs can reason over such spatial data, we extend the benchmark with DSR-Bench-spatial, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space.
-These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics.
-Given task complexity, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps for Section 4.2 and Section 4.3.
 Performance degrades as input dimensionality increases.
-As shown in Table 5, accuracy consistently drops for both models as dimensionality increases.
-Higher-dimensional data challenges models to reason over more complex distance metrics and partitions, limiting their effectiveness in spatial tasks.
-For instance, K-D trees are commonly used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision [40].
-However, a notable exception is the Geometric Graph, where 2D data performs best, likely because it is more commonly found in textbook examples used as training data.
+As shown in Table˜5, accuracy declines for both models as dimensionality increases.
+Higher-dimensional data challenges models with more complex computation over distance metrics and partitions, limiting their effectiveness in spatial tasks.
+For instance, K-D trees are widely used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision (Silpa-Anan & Hartley, 2008).
+Interestingly, 2D outperforms 1D in Geometric Graph, likely due to its more common presence in training data from textbooks.
 | Distribution | GPT-4.1 | o4-mini |
@@
 We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data.
-Specifically, we test K-D tree construction tasks using three non-uniform distributions from scikit-learn [37]: circles, moons, and blobs (examples shown in Figure 3, Section A.6).
-As shown in Table 6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in training data.
+We test K-D tree construction tasks using three non-uniform distributions from scikit-learn (Pedregosa et al., 2011): circles, moons, and blobs (illustration in Figure˜3, Section˜A.7).
+As shown in Table˜6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in the training data.
 Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning.
 In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts.
-4.3 Can LLMs reason structurally on natural language?
-While the previous sections evaluated LLMs on canonical data structures, real-world use-cases are often expressed in natural language.
-To bridge this gap, we extend the benchmark with DSR-Bench-natural, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
+A more in-depth inspection of errors and discussion on root causes can be found in Section˜A.7.
+4.3 Can LLMs reason structurally on natural language tasks?
+While the previous sections evaluated LLMs on canonical data structures, real-world use cases are often expressed in natural language.
+To bridge this gap, we extend the benchmark with the natural suite, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
 | Model | Task | Formal | Natural |
@@
 | Graph | 0.84 | 0.67 |
-We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling).
-Synthetic data follows the same distributions as Section 2.2, with realistic substitutions (e.g., names for integers).
-Each scenario was written by humans, paraphrased by GPT-4o.
+We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling game).
+Synthetic data follows the same distributions as Section˜3.3, with realistic substitutions (e.g., names for integers).
+Each scenario was written by humans and paraphrased by GPT-4o.
 All prompts were reviewed by three annotators for clarity and unambiguity.
-Details of prompt generation can be found in Section A.7.
+Details are supplemented in Section˜A.8.
 LLMs struggle when shifting from formal to natural language.
-As shown in Table 7, performance drops when tasks are described in natural language from formal descriptions, despite identical operation distributions.
-This suggests that even reasoning models struggle to apply their abilities in language-rich, real-world contexts.
+As shown in Table˜7, performance drops when tasks are described in natural language compared to formal descriptions, despite identical problem distribution.
+The higher accuracy on formal descriptors may stem from training on textbook-style patterns, where integers and explicit syntax are common.
+This observation suggests that even reasoning models struggle to apply reasoning in language-rich, real-world contexts.
 Bridging this gap is crucial for reliable deployment and presents a key direction for future research.
-Furthermore, the higher accuracy on formal descriptors may stem from LLMs’ exposure to textbook-style patterns, where integers and explicit syntax are common.
-5 Related Works
-LLM reasoning benchmarks
-Many benchmarks have been proposed to evaluate the capabilities and limitations of LLMs (see Section A.1 for a detailed review).
-Reasoning-focused benchmarks have primarily been domain-specific, targeting areas such as math [5, 28, 25], coding challenges [50, 16, 3, 58], STEM question answering [15], and logic puzzles [50, 11].
-These often require long, complex responses involving intertwined reasoning steps, and rely on human or LLM-based judgment for evaluation [4, 10, 54].
-In contrast, we take a more foundational approach: testing whether LLMs can understand and manipulate data relationships, which is a core capability that underlies reasoning in math and programming, isolating structural reasoning from other complexities.
-Algorithmic reasoning with LLMs
-Prior work on algorithmic reasoning has primarily focused on arithmetic tasks and length generalization in transformers [59, 60, 21], or on specific graph problems such as cycle detection and connectivity [48, 8].
-These efforts offer limited coverage of broader data structure reasoning.
-One closely related work, CLRS-Text [32], evaluates whether LLMs can simulate 30 algorithms from the CLRS benchmark suite [47, 6] via textual inputs.
-Rather than testing whether LLMs can replicate specific algorithms, we evaluate their ability to reason over data structures, which are core abstractions and prerequisites for algorithmic execution.
-This offers a more general and interpretable assessment of algorithmic reasoning, revealing not only whether models succeed, but also which types of structural reasoning they lack.
-6 Conclusion, limitations, and future work
-We introduce DSR-Bench, the first benchmark for evaluating LLMs’ structural reasoning through canonical data structures.
-DSR-Bench enables fine-grained, controlled, and objective evaluation using synthetic data and an automated scoring pipeline.
-Benchmarking nine state-of-the-art LLMs reveals key limitations across both instruction-tuned and reasoning models.
-Notably, DSR-Bench-challenge remains difficult even for the strongest models.
-In addition, we propose two extensions, DSR-Bench-spatial and DSR-Bench-natural, to evaluate spatial reasoning and natural language understanding of data structures.
-These results uncover further gaps, suggesting current LLMs are not yet ready for structure-dependent, real-world tasks.
-Limitation and future work
-While our scoring system compares the final state of the data structure, additional insights could be gained by evaluating models’ intermediate thought traces.
-However, such traces are unavailable from some providers and can be costly to retrieve.
-As interfaces grow more transparent, we hope to incorporate such signals to enrich error analysis and better understand model reasoning.
-Additionally, while DSR-Bench covers a wide range of structures, expanding it to cover emerging representations, such as graph-probabilistic hybrid structures, would further broaden its scope.
-Our benchmark is modular and extensible, and we welcome community contributions to advance these directions and help develop an even more comprehensive evaluation framework.
-Acknowledgments and Disclosure of Funding
+4.4 Can LLMs reason structurally with code generation?
+As motivated in Section˜1 and Section˜2, our benchmark targets general reasoning independent of code execution or tool use.
+Nonetheless, to assess whether code generation provides any benefit, we run ablations on six models using the code suite across three modes: (i) CodeOnly, where models generate Python code executed by an external interpreter; (ii) CodeEnforce, where models must write code and reason through its execution internally without relying on an interpreter; and (iii) CodeMaybe, similar to CodeEnforce but makes code generation optional.
+Full details on the experimental setup and results are provided in Section˜A.9.
+| Mode | GPT- 4.1 | o4- mini | Gemini- 2.0-Flash | Gemini- 2.5-Pro | Claude-3. 5-Sonnet | Claude-3. 7-Sonnet |
+| Default | 0.40 | 0.73 | 0.38 | 0.55 | 0.44 | 0.55 |
+| CodeMaybe | 0.38 | 0.76 | 0.41 | 0.55 | 0.43 | 0.57 |
+| CodeEnforce | 0.38 | 0.75 | 0.41 | 0.53 | 0.42 | 0.55 |
+| CodeOnly | 0.95 | 0.82 | 0.44 | 0.57 | 0.74 | 0.87 |
+| o4- mini | Gemini- 2.5-pro | Claude-3. 7-Sonnet | |
+| Geom Graph | 0.99 | 0.98 | 0.96 |
+| DAWG | 0.56 | 0.90 | 0.89 |
+| Graph-Natural | 0.69 | 0.56 | 0.23 |
+Models cannot reason over generated code.
+In Table˜9, performance in CodeMaybe and CodeEnforce matches the default setup, showing that writing code offers little benefit over natural language reasoning when models must internally simulate execution.
+This reinforces our central claim: LLMs remain limited in their ability to perform structural reasoning, even when guided by their own code.
+Code helps only with standard tasks and fails on natural language ones.
+As shown in Table˜9, with an external interpreter in CodeOnly, models perform well on Geom Graph, a standard structure in computer graphics with widely available implementations.
+In contrast, they struggle on the less familiar DAWG, where custom constraints enforce unambiguous outputs, suggesting reliance on memorized solutions rather than genuine reasoning.
+Performance drops further on Graph-Natural, where models default to brittle pattern matching (e.g., rigidly mapping “A space tunnel links A and B” to “G.add_edge(A, B)”) but failing to understand paraphrases like “Couriers frequently travel the tunnel connecting A to B.” These results highlight the fragility of structural reasoning under natural language ambiguity, even with code generation and external execution.
+5 Discussion and conclusion
+Can LLMs reason structurally?
+Through DSR-Bench, we provide a systematic answer: not yet.
+Instruction-tuned models struggle with multi-attribute reasoning (e.g., database indexing) and multi-hop reasoning (e.g., trip planning), while reasoning models achieve only 0.498 accuracy on complex structures and can ignore user-defined constraints.
+These limitations highlight the need for architectures that support precise function computation, memory mechanisms, and the flexibility to adapt to personalized requests.
+Evaluations on high-dimensional data (spatial) and natural language scenarios (natural) further reveal gaps between current reasoning capabilities and real-world readiness.
+Code generation modes (code) show that models cannot reliably reason over their own code, often reverting to memorized patterns or brittle mappings even with external execution.
+DSR-Bench provides a systematic framework for evaluating algorithmic reasoning through the lens of structural reasoning.
+It provides the community with a powerful diagnostic tool: researchers can pinpoint failure modes, test targeted improvements, and measure progress on specific relationship types, paving the way for algorithm-centric model design as an alternative to scaling alone.
+It also raises new questions: Can LLMs dynamically choose reasoning strategies?
+Where and why do errors arise in intermediate steps?
+We invite the community to use DSR-Bench to explore these directions.
+Acknowledgments
 We thank Han Xuanyuan, Connor Lawless, and members of the Vitercik Lab for their meaningful discussions and valuable feedback on the manuscript.
@@
 Appendix A Appendix
-A.1 Additional related works on LLM benchmarking
+A.1 Additional related work on LLM benchmarking
 Large language models (LLMs) have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations.
-Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding [49, 9], text generation [33, 41], reasoning [9, 46], and machine translation [53].
-Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness [42, 31, 19, 24, 57, 29].
-Specialized benchmarks have also emerged in scientific and technical domains, including mathematics [50, 25], programming [50, 16, 3, 58], data analysis [43, 27, 50, 44, 17], and medicine [23, 26, 30, 52, 18].
-While several studies examine how LLMs convert unstructured inputs into tabular or relational formats [44, 17], our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs?
+Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding (Wang et al., 2024; Feng et al., 2024), text generation (Maynez et al., 2023; Singh et al., 2024), reasoning (Feng et al., 2024; Valmeekam et al., 2023), and machine translation (Yao et al., 2023).
+Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness (Siska et al., 2024; Luo et al., 2024; Koo et al., 2023; Li et al., 2024b; Zhang et al., 2024b; Liu et al., 2023c).
+Specialized benchmarks have also emerged in scientific and technical domains, including mathematics (White et al., 2025; Liu et al., 2024a), programming (White et al., 2025; Jain et al., 2025; Chen et al., 2021; Zheng et al., 2023), data analysis (Sui et al., 2024; Liu et al., 2023b; White et al., 2025; Tang et al., 2023; Jiang et al., 2023), and medicine (Li et al., 2024a; Liu et al., 2023a; Longwell et al., 2024; Xu et al., 2024b; Kanithi et al., 2024).
+While several studies examine how LLMs convert unstructured inputs into tabular or relational formats (Tang et al., 2023; Jiang et al., 2023), our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs?
 To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
 A.2 Details of data structures and operations
-In this section, we list the data structures and the corresponding operations tested in Table 8.
+In this section, we list the data structures and the corresponding operations tested in Table˜10.
 We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
@@
 | Queue | FIFO (First-In, First-Out) | Compound | OS management | |
-| LRU Cache | Least-recently-used | Cache (Evict, Add) | Web broswers | |
+| LRU Cache | Least-recently-used | Cache (Evict, Add) | Web browsers | |
 | Priority Queue* | Priority ordering | Compound | Job scheduling | |
-| Associative | Hashmap | Key-value storage | Compound | Large-scaled storage |
+| Associative | Hashmap | Key-value storage | Compound | Large-scale storage |
 | Trie* | Hierarchical mapping of strings | Compound | Autocomplete | |
@@
 | K-D Tree* | Hierarchical, spatial partition | Construct, Compound | 3D graphics | |
-| K-D Heap* | Hierarchical, complete binary tree, high dimensional priority | Compound | GPU job scheduling | |
+| K-D Heap* | Hierarchical, complete binary tree, high-dimensional priority | Compound | GPU job scheduling | |
 | Network | Graph | Many-to-many relationships | Breadth-First Traversal, Depth-First Traversal | Social networks |
@@
 Stepwise
-This method expliclty adds a step attribute in the structured output, guiding the model to produce operations in a sequential and interpretable manner.
+This method explicitly adds a steps attribute in the JSON schema of Structured Output, guiding the model to produce operations in a sequential and interpretable manner.
+Below is an example schema for an array task:
+class Step(BaseModel): explanation: str output: str class ArraySchema(BaseModel): steps: list[Step] final_answer: int
 0-CoT
@@
 In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level.
-Table 9 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels.
-Table 10 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite.
-For detailed per-model results across all tasks and length levels, see Table 11 (o4-mini), Table 12 (Gemini-2.5-Pro), Table 13 (Claude-3.7-Sonnet), Table 14 (DeepSeek-R1), Table 15 (GPT-4.1), Table 16 (Gemini-2.0-Flash), Table 17 (Claude-3.5-Sonnet), Table 18 (DeepSeek-V3), and Table 19 (Llama-3.3).
+Table˜11 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels.
+Table˜12 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite.
+For detailed per-model results across all tasks and length levels, see Table˜13 (GPT-5), Table˜14 (o4-mini), Table˜15 (Gemini-2.5-Pro), Table˜16 (Claude-3.7-Sonnet), Table˜17 (DeepSeek-R1), Table˜18 (GPT-4.1), Table˜19 (Gemini-2.0-Flash), Table˜20 (Claude-3.5-Sonnet), Table˜21 (DeepSeek-V3), and Table˜22 (Llama-3.3).
 | Category | DS | Length | GPT-4.1 | Gemini-2.0-Flash | Claude-3.5-Sonnet | DeepSeek-V3 | Llama-3.3 |
@@
 | Long | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
-| Category | Data Structure | Length | o4-mini | Gemini-2.5-Pro | Claude-3.7-Sonnet | DeepSeek-R1 |
-| Temporal | Priority Queue | Short | 0.89 | 0.87 | 0.70 | 0.92 |
-| Medium | 0.47 | 0.43 | 0.11 | 0.54 | ||
-| Long | 0.30 | 0.24 | 0.04 | 0.48 | ||
-| Associative | Trie | Short | 0.99 | 1.00 | 0.23 | 0.93 |
-| Medium | 0.73 | 0.82 | 0.00 | 0.50 | ||
-| Long | 0.32 | 0.59 | 0.00 | 0.05 | ||
-| Suffix Tree | Short | 0.96 | 0.74 | 0.96 | 1.00 | |
-| Medium | 0.87 | 0.92 | 0.98 | 0.98 | ||
-| Long | 0.37 | 0.89 | 0.79 | 0.90 | ||
-| Skip List | Short | 0.84 | 0.93 | 0.87 | 0.89 | |
-| Medium | 0.60 | 0.64 | 0.76 | 0.63 | ||
-| Long | 0.41 | 0.61 | 0.61 | 0.54 | ||
-| Hierarchical | Heap | Short | 0.61 | 0.67 | 0.70 | 0.63 |
-| Medium | 0.72 | 0.55 | 0.37 | 0.53 | ||
-| Long | 0.71 | 0.62 | 0.13 | 0.27 | ||
-| Red Black Tree | Short | 0.92 | 0.97 | 0.69 | 0.86 | |
-| Medium | 0.67 | 0.73 | 0.11 | 0.63 | ||
-| Long | 0.37 | 0.60 | 0.12 | 0.37 | ||
-| B+ Tree | Short | 0.99 | 0.97 | 0.70 | 0.49 | |
-| Medium | 0.98 | 0.93 | 0.32 | 0.23 | ||
-| Long | 0.94 | 0.96 | 0.13 | 0.21 | ||
-| K-D Heap | Short | 0.22 | 0.07 | 0.11 | 0.23 | |
-| Medium | 0.09 | 0.08 | 0.03 | 0.08 | ||
-| Long | 0.00 | 0.00 | 0.00 | 0.01 | ||
-| K-D Tree | Short | 0.59 | 0.64 | 0.00 | 1.00 | |
-| Medium | 0.43 | 0.81 | 0.00 | 0.34 | ||
-| Long | 0.38 | 0.64 | 0.00 | 0.01 | ||
-| Network | DSU | Short | 1.00 | 1.00 | 1.00 | 1.00 |
-| Medium | 1.00 | 1.00 | 1.00 | 1.00 | ||
-| Long | 0.94 | 0.99 | 0.98 | 0.99 | ||
-| Hybrid | Bloom Filter | Short | 1.00 | 0.99 | 0.44 | 0.99 |
-| Medium | 0.77 | 1.00 | 0.03 | 0.92 | ||
-| Long | 0.07 | 0.69 | 0.00 | 0.31 | ||
-| DAWG | Short | 0.49 | 0.07 | 0.17 | 0.40 | |
-| Medium | 0.20 | 0.14 | 0.00 | 0.02 | ||
-| Long | 0.06 | 0.00 | 0.00 | 0.00 | ||
-| Geom Graph | Short | 0.37 | 0.30 | 0.07 | 0.36 | |
-| Medium | 0.10 | 0.10 | 0.00 | 0.01 | ||
-| Long | 0.00 | 0.00 | 0.00 | 0.01 |
-A.4.1 Performance of o4-mini
+| Category | Data Structure | Length | o4-mini | Gemini-2.5 Pro | Claude-3.7 Sonnet | Deepseek R1 | GPT-5 (med) |
+| Temporal | Priority Queue | Short | 0.89 | 0.89 | 0.70 | 0.92 | 0.84 |
+| Medium | 0.47 | 0.41 | 0.11 | 0.54 | 0.44 | ||
+| Long | 0.30 | 0.23 | 0.04 | 0.48 | 0.28 | ||
+| Associative | Trie | Short | 0.99 | 0.93 | 0.23 | 0.93 | 0.97 |
+| Medium | 0.73 | 0.56 | 0.00 | 0.50 | 0.94 | ||
+| Long | 0.32 | 0.37 | 0.00 | 0.05 | 0.92 | ||
+| Suffix Tree | Short | 0.96 | 0.96 | 0.96 | 1.00 | 1.00 | |
+| Medium | 0.87 | 0.89 | 0.98 | 0.98 | 1.00 | ||
+| Long | 0.37 | 0.86 | 0.79 | 0.90 | 0.95 | ||
+| Skip List | Short | 0.84 | 0.94 | 0.87 | 0.89 | 0.88 | |
+| Medium | 0.60 | 0.87 | 0.76 | 0.63 | 0.66 | ||
+| Long | 0.41 | 0.53 | 0.61 | 0.54 | 0.51 | ||
+| Hierarchical | Heap | Short | 0.61 | 0.57 | 0.70 | 0.63 | 0.78 |
+| Medium | 0.72 | 0.28 | 0.37 | 0.53 | 0.71 | ||
+| Long | 0.71 | 0.29 | 0.13 | 0.27 | 0.71 | ||
+| Red Black Tree | Short | 0.92 | 0.91 | 0.69 | 0.86 | 0.92 | |
+| Medium | 0.67 | 0.47 | 0.11 | 0.63 | 0.79 | ||
+| Long | 0.37 | 0.08 | 0.12 | 0.37 | 0.59 | ||
+| B+ Tree | Short | 0.99 | 1.00 | 0.70 | 0.49 | 0.97 | |
+| Medium | 0.98 | 0.97 | 0.32 | 0.23 | 0.99 | ||
+| Long | 0.94 | 0.94 | 0.13 | 0.21 | 0.98 | ||
+| K-D Heap | Short | 0.22 | 0.23 | 0.11 | 0.23 | 0.23 | |
+| Medium | 0.09 | 0.06 | 0.03 | 0.08 | 0.08 | ||
+| Long | 0.00 | 0.00 | 0.00 | 0.01 | 0.00 | ||
+| K-D Tree | Short | 0.59 | 0.96 | 0.00 | 1.00 | 0.97 | |
+| Medium | 0.43 | 0.64 | 0.00 | 0.34 | 0.90 | ||
+| Long | 0.38 | 0.16 | 0.00 | 0.01 | 0.67 | ||
+| Network | DSU | Short | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
+| Medium | 1.00 | 0.99 | 1.00 | 1.00 | 1.00 | ||
+| Long | 0.94 | 0.90 | 0.98 | 0.99 | 1.00 | ||
+| Hybrid | Bloom Filter | Short | 1.00 | 0.94 | 0.44 | 0.99 | 1.00 |
+| Medium | 0.77 | 0.97 | 0.03 | 0.92 | 0.84 | ||
+| Long | 0.07 | 0.66 | 0.00 | 0.31 | 0.47 | ||
+| DAWG | Short | 0.49 | 0.61 | 0.17 | 0.40 | 1.00 | |
+| Medium | 0.20 | 0.08 | 0.00 | 0.02 | 0.03 | ||
+| Long | 0.06 | 0.01 | 0.00 | 0.00 | 0.00 | ||
+| Geom Graph | Short | 0.37 | 0.19 | 0.07 | 0.36 | 0.21 | |
+| Medium | 0.10 | 0.17 | 0.00 | 0.01 | 0.08 | ||
+| Long | 0.00 | 0.02 | 0.00 | 0.01 | 0.19 |
+A.4.1 Performance of GPT-5
+| Category | Data Structure | Operation | Short | Medium | Long |
+| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
+| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Reverse | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
+| Queue | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
+| LRU Cache | Cache | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
+| Priority Queue* | Compound | 0.84 (0.02) | 0.44 (0.02) | 0.28 (0.07) | |
+| Associative | Hashmap | Compound | 1.00 (0.00) | 0.89 (0.02) | 0.71 (0.10) |
+| Trie* | Compound | 0.97 (0.03) | 0.94 (0.05) | 0.92 (0.07) | |
+| Suffix Tree* | Construct | 1.00 (0.00) | 1.00 (0.00) | 0.95 (0.02) | |
+| Skip List* | Compound | 0.88 (0.05) | 0.66 (0.08) | 0.51 (0.05) | |
+| Hierarchical | BST | Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
+| Remove | 0.98 (0.02) | 1.00 (0.00) | 0.98 (0.02) | ||
+| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Pre-Order Traversal | 0.98 (0.04) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Depth | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | ||
+| Heap* | Compound | 0.51 (0.02) | 0.81 (0.08) | 0.66 (0.12) | |
+| Heapify | 0.33 (0.00) | 0.61 (0.05) | 0.76 (0.10) | ||
+| RB Tree* | Construct | 0.93 (0.00) | 0.90 (0.03) | 0.68 (0.13) | |
+| Compound | 0.91 (0.04) | 0.67 (0.03) | 0.49 (0.04) | ||
+| B+ Tree* | Compound | 0.97 (0.00) | 0.99 (0.02) | 0.98 (0.02) | |
+| K-D Tree* | Construct | 0.97 (0.03) | 0.90 (0.00) | 0.67 (0.06) | |
+| K-D Heap* | Compound | 0.23 (0.00) | 0.08 (0.02) | 0.00 (0.00) | |
+| Network | Graph | Breadth-First Traversal | 0.92 (0.02) | 0.98 (0.02) | 0.94 (0.02) |
+| Depth-First Traversal | 0.93 (0.03) | 1.00 (0.00) | 0.96 (0.02) | ||
+| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
+| Geom Graph* | Construct | 0.21 (0.10) | 0.08 (0.04) | 0.19 (0.13) | |
+| Hybrid | Bloom Filter* | Construct | 1.00 (0.00) | 0.84 (0.07) | 0.47 (0.00) |
+| DAWG* | Compound | 1.00 (0.00) | 0.03 (0.03) | 0.00 (0.00) |
+A.4.2 Performance of o4-mini
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.84 (0.02) | 0.60 (0.03) | 0.41 (0.02) | |
-| Associative | BST | Insert | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) |
+| Hierarchical | BST | Insert | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) |
 | Remove | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.03) | ||
@@
 | Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
-| Depth | 0.01 (0.02) | 0.03 (0.03) | 0.00 (0.00) | ||
+| Depth | 1.00 (0.00) | 0.99 (0.02) | 1.00 (0.00) | ||
 | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | ||
@@
 | DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.94 (0.04) | |
-| Geom Graph* | Construct | 0.37 (0.00) | 0.10 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 1.00 (0.00) | 0.77 (0.09) | 0.07 (0.00) |
+| Geom Graph* | Construct | 0.83 (0.07) | 0.13 (0.00) | 0.01 (0.02) | |
+| Hybrid | Bloom Filter* | Compound | 1.00 (0.00) | 0.77 (0.09) | 0.07 (0.00) |
 | DAWG* | Compound | 0.49 (0.10) | 0.20 (0.09) | 0.06 (0.05) |
-A.4.2 Performance of Gemini-2.5-Pro
+A.4.3 Performance of Gemini-2.5-Pro
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
-| Insert | 1.00 (0.00) | 0.98 (0.02) | 1.00 (0.00) | ||
-| Reverse | 1.00 (0.00) | 0.98 (0.02) | 1.00 (0.00) | ||
+| Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Reverse | 1.00 (0.00) | 0.98 (0.02) | 0.99 (0.02) | ||
 | Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
@@
 | LRU Cache | Cache | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
-| Priority Queue* | Compound | 0.87 (0.03) | 0.43 (0.03) | 0.24 (0.05) | |
-| Associative | Hashmap | Compound | 0.90 (0.06) | 0.51 (0.04) | 0.50 (0.07) |
-| Trie* | Compound | 1.00 (0.00) | 0.82 (0.05) | 0.59 (0.08) | |
-| Suffix Tree* | Construct | 0.74 (0.08) | 0.92 (0.08) | 0.89 (0.02) | |
-| Skip List* | Compound | 0.93 (0.03) | 0.64 (0.05) | 0.61 (0.02) | |
-| Associative | BST | Insert | 0.99 (0.02) | 0.94 (0.05) | 0.96 (0.02) |
-| Remove | 0.99 (0.02) | 0.84 (0.05) | 0.90 (0.03) | ||
-| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
+| Priority Queue* | Compound | 0.89 (0.02) | 0.41 (0.02) | 0.23 (0.05) | |
+| Associative | Hashmap | Compound | 0.58 (0.02) | 0.16 (0.07) | 0.11 (0.04) |
+| Trie* | Compound | 0.93 (0.03) | 0.56 (0.02) | 0.37 (0.03) | |
+| Suffix Tree* | Construct | 0.96 (0.04) | 0.89 (0.02) | 0.86 (0.08) | |
+| Skip List* | Compound | 0.94 (0.02) | 0.87 (0.06) | 0.53 (0.03) | |
+| Hierarchical | BST | Insert | 1.00 (0.00) | 0.94 (0.02) | 0.94 (0.05) |
+| Remove | 0.87 (0.06) | 0.86 (0.07) | 0.89 (0.02) | ||
+| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
 | Pre-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
-| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
-| Depth | 0.22 (0.02) | 0.17 (0.07) | 0.18 (0.04) | ||
-| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
-| Heap* | Compound | 0.91 (0.04) | 0.72 (0.11) | 0.71 (0.05) | |
-| Heapify | 0.43 (0.21) | 0.37 (0.03) | 0.53 (0.09) | ||
-| RB Tree* | Construct | 0.83 (0.05) | 0.60 (0.03) | 0.28 (0.07) | |
-| Compound | 0.91 (0.02) | 0.47 (0.06) | 0.18 (0.08) | ||
-| B+ Tree* | Compound | 0.97 (0.06) | 0.93 (0.03) | 0.96 (0.05) | |
-| K-D Tree* | Construct | 0.64 (0.11) | 0.81 (0.05) | 0.67 (0.12) | |
-| K-D Heap* | Compound | 0.07 (0.00) | 0.08 (0.02) | 0.00 (0.00) | |
-| Network | Graph | Breadth-First Traversal | 0.98 (0.04) | 0.99 (0.02) | 0.86 (0.12) |
-| Depth-First Traversal | 0.98 (0.02) | 0.94 (0.04) | 0.86 (0.05) | ||
-| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | |
-| Geom Graph* | Construct | 0.30 (0.03) | 0.10 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.99 (0.02) | 1.00 (0.00) | 0.69 (0.02) |
-| DAWG* | Compound | 0.07 (0.02) | 0.14 (0.08) | 0.00 (0.00) |
-A.4.3 Performance of Claude-3.7-Sonnet
+| Post-Order Traversal | 1.00 (0.00) | 0.99 (0.02) | 0.94 (0.04) | ||
+| Depth | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
+| Compound | 0.99 (0.02) | 0.99 (0.02) | 0.98 (0.02) | ||
+| Heap* | Compound | 0.78 (0.04) | 0.49 (0.10) | 0.53 (0.12) | |
+| Heapify | 0.36 (0.02) | 0.06 (0.05) | 0.04 (0.04) | ||
+| RB Tree* | Construct | 0.91 (0.07) | 0.53 (0.09) | 0.03 (0.03) | |
+| Compound | 0.91 (0.02) | 0.41 (0.07) | 0.13 (0.03) | ||
+| B+ Tree* | Compound | 1.00 (0.00) | 0.97 (0.03) | 0.94 (0.06) | |
+| K-D Tree* | Construct | 0.96 (0.02) | 0.64 (0.13) | 0.16 (0.02) | |
+| K-D Heap* | Compound | 0.23 (0.00) | 0.06 (0.02) | 0.00 (0.00) | |
+| Network | Graph | Breadth-First Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.74 (0.02) |
+| Depth-First Traversal | 1.00 (0.00) | 0.81 (0.05) | 0.14 (0.02) | ||
+| DSU* | Compound | 1.00 (0.00) | 0.99 (0.02) | 0.90 (0.05) | |
+| Geom Graph* | Construct | 0.19 (0.02) | 0.17 (0.06) | 0.02 (0.02) | |
+| Hybrid | Bloom Filter* | Construct | 0.94 (0.02) | 0.97 (0.00) | 0.66 (0.02) |
+| DAWG* | Compound | 0.61 (0.05) | 0.08 (0.07) | 0.01 (0.02) |
+A.4.4 Performance of Claude-3.7-Sonnet
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.87 (0.06) | 0.76 (0.05) | 0.61 (0.11) | |
-| Associative | BST | Insert | 0.96 (0.04) | 0.98 (0.04) | 0.79 (0.02) |
+| Hierarchical | BST | Insert | 0.96 (0.04) | 0.98 (0.04) | 0.79 (0.02) |
 | Remove | 0.94 (0.02) | 0.91 (0.02) | 0.92 (0.02) | ||
@@
 | Post-Order Traversal | 1.00 (0.00) | 0.74 (0.02) | 0.93 (0.03) | ||
-| Depth | 0.24 (0.11) | 0.13 (0.03) | 0.02 (0.04) | ||
+| Depth | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | ||
 | Compound | 0.90 (0.03) | 0.21 (0.02) | 0.24 (0.05) | ||
@@
 | DSU* | Compound | 0.04 (0.02) | 0.12 (0.04) | 0.00 (0.00) | |
-| Geom Graph* | Construct | 0.07 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) |
+| Geom Graph* | Construct | 0.04 (0.05) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.17 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
-A.4.4 Performance of DeepSeek-R1
+A.4.5 Performance of DeepSeek-R1
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.89 (0.04) | 0.63 (0.03) | 0.54 (0.02) | |
-| Associative | BST | Insert | 1.00 (0.00) | 0.98 (0.02) | 0.90 (0.03) |
+| Hierarchical | BST | Insert | 1.00 (0.00) | 0.98 (0.02) | 0.90 (0.03) |
 | Remove | 0.98 (0.02) | 0.93 (0.03) | 0.88 (0.05) | ||
@@
 | Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.97 (0.00) | ||
-| Depth | 0.02 (0.04) | 0.06 (0.04) | 0.06 (0.05) | ||
+| Depth | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
 | Compound | 0.97 (0.03) | 0.84 (0.08) | 0.65 (0.07) | ||
@@
 | DSU* | Compound | 0.64 (0.56) | 0.92 (0.04) | 0.83 (0.07) | |
-| Geom Graph* | Construct | 0.36 (0.02) | 0.01 (0.02) | 0.01 (0.02) | |
-| Network | Bloom Filter* | Compound | 0.99 (0.02) | 0.92 (0.02) | 0.31 (0.02) |
+| Geom Graph* | Construct | 0.99 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.99 (0.02) | 0.92 (0.02) | 0.31 (0.02) |
 | DAWG* | Compound | 0.40 (0.12) | 0.02 (0.02) | 0.00 (0.00) |
-A.4.5 Performance of GPT-4.1
+A.4.6 Performance of GPT-4.1
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.21 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Associative | BST | Insert | 0.79 (0.04) | 0.50 (0.03) | 0.14 (0.02) |
+| Hierarchical | BST | Insert | 0.79 (0.04) | 0.50 (0.03) | 0.14 (0.02) |
 | Remove | 0.78 (0.04) | 0.58 (0.02) | 0.36 (0.04) | ||
@@
 | Post-Order Traversal | 0.82 (0.02) | 0.51 (0.04) | 0.23 (0.06) | ||
-| Depth | 0.66 (0.05) | 0.08 (0.05) | 0.03 (0.00) | ||
+| Depth | 0.30 (0.04) | 0.07 (0.06) | 0.03 (0.03) | ||
 | Compound | 0.69 (0.02) | 0.26 (0.02) | 0.03 (0.00) | ||
@@
 | DSU* | Compound | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Geom Graph* | Construct | 0.20 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
+| Geom Graph* | Construct | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.16 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
-A.4.6 Performance of Gemini-2.0-Flash
+A.4.7 Performance of Gemini-2.0-Flash
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.16 (0.02) | 0.00 (0.00) | 0.03 (0.00) | |
-| Associative | BST | Insert | 0.31 (0.02) | 0.27 (0.03) | 0.06 (0.04) |
+| Hierarchical | BST | Insert | 0.31 (0.02) | 0.27 (0.03) | 0.06 (0.04) |
 | Remove | 0.63 (0.09) | 0.33 (0.03) | 0.13 (0.03) | ||
@@
 | Post-Order Traversal | 0.63 (0.00) | 0.17 (0.00) | 0.10 (0.00) | ||
-| Depth | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) | ||
+| Depth | 0.13 (0.09) | 0.03 (0.00) | 0.00 (0.00) | ||
 | Compound | 0.51 (0.05) | 0.12 (0.04) | 0.10 (0.00) | ||
@@
 | DSU* | Compound | 0.01 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Geom Graph* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
+| Geom Graph* | Construct | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.18 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
-A.4.7 Performance of Claude-3.5-Sonnet
+A.4.8 Performance of Claude-3.5-Sonnet
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.77 (0.06) | 0.30 (0.07) | 0.20 (0.07) | |
-| Associative | BST | Insert | 0.80 (0.06) | 0.50 (0.06) | 0.51 (0.05) |
+| Hierarchical | BST | Insert | 0.80 (0.06) | 0.50 (0.06) | 0.51 (0.05) |
 | Remove | 0.96 (0.04) | 0.87 (0.00) | 0.77 (0.00) | ||
@@
 | Post-Order Traversal | 1.00 (0.00) | 0.69 (0.08) | 0.54 (0.02) | ||
-| Depth | 0.76 (0.07) | 0.34 (0.10) | 0.16 (0.13) | ||
+| Depth | 1.00 (0.00) | 0.96 (0.02) | 0.78 (0.02) | ||
 | Compound | 0.77 (0.07) | 0.28 (0.05) | 0.09 (0.02) | ||
@@
 | DSU* | Compound | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
-| Geom Graph* | Construct | 0.06 (0.04) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
+| Geom Graph* | Construct | 0.10 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.20 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
-A.4.8 Performance of DeepSeek-V3
+A.4.9 Performance of DeepSeek-V3
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Associative | BST | Insert | 0.93 (0.03) | 0.62 (0.02) | 0.46 (0.05) |
+| Hierarchical | BST | Insert | 0.93 (0.03) | 0.62 (0.02) | 0.46 (0.05) |
 | Remove | 0.84 (0.04) | 0.80 (0.03) | 0.66 (0.02) | ||
@@
 | Post-Order Traversal | 0.82 (0.02) | 0.53 (0.03) | 0.20 (0.03) | ||
-| Depth | 0.92 (0.02) | 0.43 (0.03) | 0.08 (0.04) | ||
+| Depth | 0.67 (0.03) | 0.24 (0.02) | 0.07 (0.03) | ||
 | Compound | 0.68 (0.02) | 0.12 (0.04) | 0.06 (0.02) | ||
@@
 | DSU* | Compound | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
-| Geom Graph* | Construct | 0.04 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
+| Geom Graph* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
+| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.17 (0.00) | 0.00 (0.00) | 0.00 (0.00) |
-A.4.9 Performance of Llama-3.3
+A.4.10 Performance of Llama-3.3
 | Category | Data Structure | Operation | Short | Medium | Long |
@@
 | Skip List* | Compound | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
-| Associative | BST | Insert | 0.37 (0.00) | 0.14 (0.02) | 0.03 (0.00) |
+| Hierarchical | BST | Insert | 0.37 (0.00) | 0.14 (0.02) | 0.03 (0.00) |
 | Remove | 0.49 (0.04) | 0.30 (0.03) | 0.14 (0.04) | ||
@@
 | Post-Order Traversal | 0.31 (0.04) | 0.04 (0.02) | 0.00 (0.00) | ||
-| Depth | 0.33 (0.03) | 0.23 (0.00) | 0.28 (0.05) | ||
+| Depth | 0.70 (0.00) | 0.38 (0.07) | 0.13 (0.06) | ||
 | Compound | 0.26 (0.02) | 0.01 (0.02) | 0.00 (0.00) | ||
@@
 | Geom Graph* | Construct | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
-| Network | Bloom Filter* | Compound | 0.00 (0.00) | 0.07 (0.00) | 0.00 (0.00) |
+| Hybrid | Bloom Filter* | Compound | 0.00 (0.00) | 0.07 (0.00) | 0.00 (0.00) |
 | DAWG* | Compound | 0.02 (0.02) | 0.00 (0.00) | 0.00 (0.00) |
@@
 This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None.
-The results are shown in Table 20 (GPT-4.1), Table 21 (Gemini-2.0-Flash), Table 22 (Claude-3.5-Sonnet), Table 23 (DeepSeek-V3), and Table 24 (Llama-3.3).
+The results are shown in Table˜23 (GPT-4.1), Table˜24 (Gemini-2.0-Flash), Table˜25 (Claude-3.5-Sonnet), Table˜26 (DeepSeek-V3), and Table˜27 (Llama-3.3).
 | Data structure | Task | Stepwise | 0-CoT | CoT | 3-shot | None |
@@
 | DAWG | Compound |
-A.6 DSR-Bench-spatial supplementary materials
-Figure 3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn.
-These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section 4.2.
-A.7 DSR-Bench-natural supplementary materials
-Figure 4 illustrates our generation process for natural language prompts in DSR-Bench-natural.
-For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in A.7).
+A.6 Additional analysis on CoT prompting
+In Section˜4.1.1, we analyzed various prompting strategies with instruction-tuned models.
+We further examined the CoT method by inspecting its reasoning outputs, which offers insights we hope will benefit practitioners and researchers:
+CoT offers limited benefits for well-known tasks.
+For familiar problems that models are likely to have encountered during pretraining (such as array, queue, and Binary Search Tree), CoT prompting yields only marginal gains.
+This suggests that models already possess internalized procedures for these tasks and can execute them reliably without additional reasoning steps.
+Simpler prompts can be more cost efficient and effective.
+Without careful design, CoT can hurt performance.
+In more complex scenarios, such as the hashmap compound task, using CoT with structured JSON-style reasoning actually degraded accuracy.
+When the reasoning format was changed to natural language, performance recovered to match the baseline (0-CoT).
+This highlights that the effectiveness of CoT is highly sensitive to prompt design, and poorly chosen reasoning formats may introduce unnecessary complexity rather than aiding problem solving.
+A.7 The spatial suite supplementary materials
+Figure˜3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn.
+These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section˜4.2.
+We observe a performance drop in KD-Tree tasks when the input data is non-uniform.
+At the problem level, uniform and non-uniform KD-Tree tasks appear equally difficult for a human reasoner.
+Across 30 questions per group, duplicated indices, which trigger tie-breaking, occur at similar rates on both axes (x: 32 vs.
+33; y: 26 vs.
+28), and the total number of median operations is comparable (156 vs.
+145).
+This suggests the surface-level difficulty is well matched across distributions.
+A closer inspection of errors reveals two main causes:
+Non-uniform data often increases ambiguity near the global median.
+Non-uniform distributions often produce clusters of similar or identical coordinates near the global median, making it harder for the model to select the correct split.
+Since the root node is determined by a median over the entire dataset, this ambiguity leads to more errors.
+Supporting this, we observed 13 out of 30 correct root nodes on uniform data, but only 5 out of 30 on circle (non-uniform) data.
+Even simple tasks, such as median calculation, can be challenging for LLMs.
+On non-uniform data, models exhibited 28.6% more cases of violating the prompt’s “latter-median” rule by instead choosing the “former-median,” and 50% more instances of selecting a completely incorrect median (neither the former nor latter).
+Additionally, while no axis confusion errors occurred on uniform data, the model made three such errors on non-uniform data, mistakenly selecting the wrong axis before performing median calculation.
+A.8 The natural suite supplementary materials
+Figure˜4 illustrates our generation process for natural language prompts in DSR-Bench-natural.
+For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in Section˜A.8).
 This narrative is then paraphrased into five variants using GPT-4o.
-For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in A.7).
+For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in Section˜A.8).
 In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
@@
 We construct a scenario in which a clinic uses a BST to store patient appointments, ordered by appointment time and tie-broken by the patient’s name.
-The insert operation adds a name, time) appointment to the tree, while the delete operation corresponds to a patient canceling their appointment.
+The insert operation adds a (name, time) appointment to the tree, while the delete operation corresponds to a patient canceling their appointment.
 To retrieve all records, a pre-order traversal of the tree is performed.
@@
 The task is to navigate a starship to visit as many planets as possible using depth-first search, starting from a given planet and visiting neighbors in lexicographical order.
-A.8 Levenshtein distance: an auxiliary metric for DSR-Bench
-In addition to the binary (0/1) accuracy reported in the main text, DSR,-Bench includes an optional evaluation metric based on Levenshtein distance, which measures the minimum number of single-character insertions, deletions, or substitutions needed to transform one string into another.
+A.9 The code suite supplementary materials
+Our benchmark is designed to evaluate general reasoning in LLMs without relying on external tools such as code interpreters or formal solvers.
+The focus is on the underlying reasoning ability that drives broad problem-solving, beyond coding or domains where code is directly applicable.
+This follows recent evaluation practices, such as Gemini-Deep-Think and OpenAI’s IMO assessments (Luong & Lockhart, 2025; Wei, 2025), which deliberately disallowed code or tool use (e.g., Lean) to test end-to-end reasoning.
+By using data structures as controlled, interpretable settings for structural reasoning—a core component of general reasoning—DSR-Bench differs from code-synthesis benchmarks that emphasise syntax or API knowledge and are often vulnerable to data leakage.
+To additionally probe how code generation may support structural reasoning, we conducted an ablation across six models, including reasoning models with competitive performance on code-synthesis benchmarks.
+We cover seven data structures of varying difficulty (Array, Queue, Hashmap, Heap, DAWG, Geom Graph, Graph-Natural) and average results over three runs.
+Three code-generation modes were tested, with prompts shown below:
+CodeOnly: “Your answer should be a Python function ‘def solution()’ that takes no inputs, solves the problem, and returns the final solution in the expected format. Output the code only.”
+CodeEnforce: “You should write code to solve the problem, then reason through its execution to explain what the output would be. Your final answer should be the output itself.”
+CodeMaybe: “You can write code to help solve the problem, or solve it directly.
+If you use code, reason through its execution to explain what the output would be.
+Your final answer should be the output itself.”
+| GPT-4.1 | o4-mini | |||||||
+| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | None | CodeMaybe | CodeEnforce | CodeOnly |
+| Array | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
+| Queue | 0.82 | 0.84 | 0.89 | 0.98 | 1.00 | 1.00 | 1.00 | 1.00 |
+| Hashmap | 0.19 | 0.11 | 0.09 | 1.00 | 0.89 | 0.89 | 0.86 | 0.94 |
+| Heap | 0.58 | 0.53 | 0.53 | 1.00 | 0.44 | 0.73 | 0.70 | 0.53 |
+| DAWG | 0.16 | 0.17 | 0.11 | 0.90 | 0.49 | 0.46 | 0.30 | 0.56 |
+| Geom Graph | 0.03 | 0.02 | 0.04 | 0.93 | 0.83 | 0.98 | 0.97 | 0.99 |
+| Graph-Natural | 0.01 | 0.00 | 0.00 | 0.86 | 0.43 | 0.26 | 0.39 | 0.69 |
+| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | Accuracy | Frequency |
+| Array | 1.00 | 1.00 | 1.00 | 0.01 | 1.00 | 0.01 |
+| Queue | 0.87 | 0.86 | 0.87 | 0.13 | 0.93 | 0.14 |
+| Hashmap | 0.28 | 0.49 | 0.49 | 0.62 | 0.93 | 0.67 |
+| Heap | 0.23 | 0.28 | 0.34 | 0.83 | 1.00 | 0.83 |
+| DAWG | 0.18 | 0.20 | 0.17 | 0.00 | 0.00 | 0.01 |
+| Geom Graph | 0.07 | 0.03 | 0.03 | 0.68 | 0.76 | 0.90 |
+| Graph-Natural | 0.00 | 0.00 | 0.00 | 0.82 | 0.82 | 1.00 |
+| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | Accuracy | Frequency |
+| Array | 1.00 | 1.00 | 1.00 | 0.83 | 1.00 | 0.83 |
+| Queue | 1.00 | 1.00 | 1.00 | 0.99 | 1.00 | 0.99 |
+| Hashmap | 0.58 | 0.17 | 0.13 | 0.57 | 0.95 | 0.60 |
+| Heap | 0.36 | 0.37 | 0.50 | 0.37 | 0.42 | 0.87 |
+| DAWG | 0.61 | 0.47 | 0.45 | 0.33 | 0.90 | 0.37 |
+| Geom Graph | 0.19 | 0.77 | 0.53 | 0.64 | 0.98 | 0.65 |
+| Graph-Natural | 0.12 | 0.10 | 0.09 | 0.24 | 0.56 | 0.43 |
+| Claude-3.5-Sonnet | Claude-3.7-Sonnet | |||||||
+| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | None | CodeMaybe | CodeEnforce | CodeOnly |
+| Array | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
+| Queue | 0.87 | 0.86 | 0.84 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
+| Hashmap | 0.37 | 0.27 | 0.27 | 1.00 | 0.71 | 0.73 | 0.65 | 1.00 |
+| Heap | 0.53 | 0.63 | 0.63 | 0.93 | 0.89 | 0.85 | 0.82 | 0.98 |
+| DAWG | 0.20 | 0.17 | 0.13 | 0.20 | 0.17 | 0.18 | 0.15 | 0.89 |
+| Geom Graph | 0.10 | 0.09 | 0.10 | 0.98 | 0.04 | 0.23 | 0.20 | 0.96 |
+| Graph-Natural | 0.00 | 0.00 | 0.00 | 0.10 | 0.01 | 0.02 | 0.01 | 0.23 |
+While GPT-4.1 and o4-mini generally perform well, we observe that Claude models also perform competitively, particularly in spatial data structure tasks such as Geom Graph.
+The only exception is Graph-Natural, suggesting that Claude models struggle with handling natural language ambiguity.
+On the other hand, Gemini models tend to return final answers directly rather than code in CodeOnly mode, despite explicit instructions (“Your answer should be a Python function… Output the code only.”).
+This indicates weaker instruction-following in code generation compared to other models and contributes to their low performance.
+We investigated this further and presented a new table showing (i) task accuracy when code is written (Code Accuracy) and (ii) the code-writing frequency of Gemini models (Code Frequency).
+We observe that Gemini models also perform on par when code is written.
+Models cannot reason over the code they write.
+Performance in CodeMaybe and CodeEnforce is on par with our original setup, indicating that writing code does not improve reasoning when models must internally simulate it.
+This reinforces our central claim: LLMs still struggle with structural reasoning, even when guided by their own code.
+Code helps when tasks align with memorized patterns.
+In CodeOnly, models perform well on Geom Graph (k-dimensional graphs embedded in geometric space, which is a standard data structure widely used in computer graphics), whose code implementation is more available online.
+However, they struggle with the less familiar DAWG” (directed acyclic word graph), where we define it with customized constraints to enforce a unique output.
+This suggests that performance may reflect memorization rather than true reasoning.
+CodeOnly falters on natural language task variants.
+In Graph-Natural, models rely on brittle pattern matching (e.g., mapping “A space tunnel links planet1 and planet2” to “G.add_edge(planet1, planet2)”), but often fail to cover all phrasing variations or misinterpret descriptions (e.g., missing “Couriers frequently travel the tunnel that connects planet1 to planet2”).
+This highlights a key limitation that models still struggle to apply structural reasoning to ambiguous natural language scenarios, even with external tools.
+Ablation with ChatGPT web UI.
+Our benchmark was evaluated via APIs.
+In contrast, the ChatGPT web UI has a built-in code interpreter that executes code automatically when needed.
+We tested o4-mini on ChatGPT with DAWG+CodeMaybe.
+The model scored 0.70, generating and executing code in 20 out of 30 cases.
+We observed that the model typically uses code for complex algorithmic components, while relying on natural language reasoning for the rest.
+A.10 Levenshtein distance: an auxiliary metric for DSR-Bench
+In addition to the binary (0/1) accuracy reported in the main text, DSR-Bench includes an optional evaluation metric based on Levenshtein distance, which measures the minimum number of single-character insertions, deletions, or substitutions needed to transform one string into another.
 As a continuous metric, it captures degrees of error that binary accuracy flattens.
-For instance, given the correct output [1,,3,,6], the prediction [3,,1,,6] is clearly closer than [0,,0,,0], and Levenshtein distance reflects that nuance.
+For instance, given the correct output [1,3,6], the prediction [3,1,6] is clearly closer than [0,0,0], and Levenshtein distance reflects that nuance.
 However, this granularity can also blur important semantic distinctions.
@@
 When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein.
-For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table 25).
+For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table˜32).
 Output length further complicates cross-task comparison.
@@
 The implementation remains publicly available, as the metric can still offer a useful secondary perspective, particularly when comparing models on the same task and output length.
+A.11 Failure rates of JSON parsing via Structured Output
+| Model | Array | Priority Queue | Hashmap | RB Tree | Geom Graph | Bloom Filter |
+| Llama3.3 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| GPT-4.1 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| DeepSeek-Chat | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| DeepSeek-R1 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| o4-mini | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| Claude-3.5-Sonnet | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| Claude-3.7-Sonnet | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| Gemini-2.0-Flash | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+| Gemini-2.5-Pro | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
+A.12 Ablation on paraphrased prompt templates
+In the natural suite, we use 5 paraphrased templates for problem statements and 30 for operations (e.g., “A space tunnel connects planet1 and planet2”) to smooth out ambiguity in natural language descriptions.
+However, in the main and challenge suites, we use formalized prompt templates to enable large-scale evaluation.
+We conducted an ablation to test how prompt template variations affect performance.
+Prompting GPT-4o with “Paraphrase the following description,” we generated five paraphrases and found that overall performance trends remained consistent.
+| o4-mini | GPT-4.1 | |||
+| Task | Default | Paraphrased | Default | Paraphrased |
+| Priority Queue | 0.99 | 0.84 | 0.39 | 0.36 |
+| Trie Tree | 0.89 | 0.90 | 0.63 | 0.61 |
+| Geom Graph | 0.37 | 0.37 | 0.20 | 0.14 |
+A.13 Preliminary study on strategy adaptation with DSR-Bench
+In Section˜5, we note a future direction is to study whether models can select appropriate data structures given task requirements and switch dynamically.
+DSR-Bench already implicitly evaluates this adaptivity by not predefining strategies in prompts, requiring models to adjust their reasoning to the task.
+This emerges in (i) contrasting but related structure pairs (e.g., BFS/DFS, queue/priority queue), (ii) length generalization that demands adaptive complexity management, (iii) distribution shifts (Section˜4.2) requiring robustness across input patterns, and (iv) natural language scenarios that demand transferring strategies across contexts (Section˜4.3).
+Beyond implicit evaluation, DSR-Bench’s modular design—data generation, prompt templates, automated evaluation, and schema-based verification—readily supports dynamic strategy switching.
+A preliminary study is shown in Table˜36 and Table˜35.
+We mask structure names or add constraints to DSR-Bench’s prompts, and test whether models can, (i) select BFS vs DFS for graph tasks, or (ii) choose between K-D tree and array for nearest-neighbor queries under varying complexities.
+This extension underscores DSR-Bench’s broader value: it serves as a fundamental, extensible framework for probing not only structural reasoning but also more advanced reasoning abilities.
+| Task | GPT-4.1 | o4-mini |
+| Shortest Path | 9.5/10 (BFS) | 10/10 (BFS) |
+| Cycle Detection | 10/10 (DFS) | 10/10 (DFS) |
+| Connectivity Detection | 10/10 (Both equal; 8 BFS, 2 DFS) | 10/10 (Both equal; 6 BFS, 4 DFS) |
+| Task | GPT-4.1 | o4-mini |
+| 1 query, n=20–30 | 9/10 (array) | 10/10 (array) |
+| 1 query, n=40–60 | 6/10 (array) | 10/10 (array) |
+| n queries, n=5–10 | 10/10 (array) | 10/10 (array) |
+| n queries, n=20–30 | 10/10 (kd-tree) | 10/10 (kd-tree) |